### v1

# Continual Release Moment Estimation with Differential Privacy

Abstract
We propose Joint Moment Estimation (JME), a method for continually and privately estimating both the first and second moments of data with reduced noise compared to naive approaches. JME uses the matrix mechanism and joint sensitivity analysis to allow the second-moment estimation with no additional privacy cost, thereby improving accuracy while maintaining privacy. We demonstrate JME’s effectiveness in two applications: estimating the running mean and covariance matrix for Gaussian density estimation and model training with DP-Adam on CIFAR-10.
Machine Learning, ICML
1 Introduction
Estimating the first and second moments of data is a fundamental step in many machine learning algorithms, ranging from foundational methods, such as linear regression, principal component analysis, or Gaussian model fitting, to state-of-the-art neural network components, such as BatchNorm (Ioffe & Szegedy, 2015), and optimizers, such as Adam (Diederik & Ba, 2015). Often, these estimates must be computed and updated continuously, called the continual release setting (Dwork et al., 2010). For example, this is the case when the data arrives sequentially and intermediate results are needed without delay, such as for sequential optimization algorithms, such as Adam, or for real-time systems in healthcare, finance, or online recommendation systems. In such scenarios, ensuring the privacy of sensitive data, such as user information or medical records, is essential.
Differential Privacy (DP) is a widely established formal notion of privacy that ensures that the inclusion or exclusion of any single individual’s data has a limited impact on the output of an algorithm. Technically, DP masks sensitive information by the addition of suitably scaled noise, thereby creating a trade-off between privacy (formalized as a privacy budget) and utility (measured by the expected accuracy of the estimates). This trade-off becomes particularly challenging when more than one quantity is meant to be estimated from the same private data, as is the case when estimating multiple data moments. Done naively, the privacy budget has to be split between the estimates, resulting in more noise and lower accuracy for both of them. In this work, we introduce a new method, Joint Moment Estimation (JME), that is able to privately estimate the first and second moments of vector-valued data without suffering from this shortcoming. Algorithmically, JME relies on the recent matrix factorization (MF) mechanism (Li et al., 2015) for continual release DP to individually compute the first and the second moments of the data, thereby making it flexible to accommodate a variety of settings. For example, besides the standard uniform sum or weighted average across data items, exponentially weighted averages or sliding-window estimates are readily possible. The key innovation of JME lies in our theoretical analysis of its properties. By jointly analyzing the sensitivity of the otherwise independent estimation processes, in combination with considering a carefully calibrated trade-off between them, we show that one can privately estimate the second-moment matrix without having to increase the amount of noise required to keep the first moment private. In this sense, we obtain privacy of the second moment for free. JME is practical and easily implemented using standard programming languages and toolboxes. We demonstrate this by showcasing two applications, one classical and one modern. First, we use JME for continual density estimation using a multivariate Gaussian model, e.g. we estimate the running mean and covariance matrix of a sequence of vector-valued observations. Our experiments confirm that JME achieves a lower Frobenius norm error for the covariance matrix in high-privacy regimes as well as a better fit to the true distribution as measured by the Kullback-Leibler divergence. Second, we integrate JME into the Adam optimizer, which is widely used in deep learning. Here, as well, we observe that JME achieves better optimization accuracy than baseline methods in high-privacy and small-batch-size regimes.
method symbol coefficients
prefix sum $E_{1}$ $\mathbbm{1}{\{i\leq t\}}$
exponential (weight $\beta$) $E_{\beta}$ $\beta^{t-i}\mathbbm{1}\{i\leq t\}$
standard average $V$ $\frac{1}{t}\mathbbm{1}\{i\leq t\}$
sliding window (size $k$) $W_{k}$ $\frac{1}{k}\mathbbm{1}\{t-k\,<\,i\,\leq\,t\}$
Table 1: Common workload matrices for moment estimation in (1).
2 Background
We study the problem of continually estimating the pair of weighted sums of the first and second moments in a differentially private manner. Namely, consider a sequence of $d$-dimensional vectors $x_{1},\dots,x_{n}\in\mathcal{X}$, where $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ for some fixed constant $\zeta>0$. At each step $t$, we aim to estimate the following pair of sums in a differentially private way:
$Y_{t}$ $=\sum\limits_{i=1}^{t}a^{t}_{i}x_{i}\in\mathbb{R}^{d}\ \ \text{and}\ \ S_{t}=\sum\limits_{i=1}^{t}b^{t}_{i}x_{i}x_{i}^{\top}\in\mathbb{R}^{d\times d},$ (1)
for arbitrary coefficients $a^{t}_{i}\in\mathbb{R}$ and $b^{t}_{i}\in\mathbb{R}$. This formulation includes many practical schemes, see Table 1.
To express this problem compactly, we rewrite it in matrix form. We collect the coefficients into lower triangular workload matrices, $A_{1}=(a_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$ and $A_{2}=(b_{i}^{t})_{1\leq i,t\leq n}\in\mathbb{R}^{n\times n}$. We stack the data as rows into a matrix $X\in\mathbb{R}^{n\times d}$. Then, all terms of the sums in (1) can then be expressed compactly [^1] For simplicity, we slightly abuse the notation and make $S$ a matrix of shape $n\times d^{2}$ instead of a tensor of size $n\times d\times d$. as $Y=A_{1}X$ and $S=A_{2}(X\operatorname{\!\bullet\!}X)$, where the second matrix consists of the data vectors stacked as rows, and $\bullet$ denotes the Face-Splitting Product or transposed Khatri-Rao product (Esteve et al., 2009), $(X\operatorname{\!\bullet\!}X)_{ij}:=(x_{i}\otimes x_{i})_{j}=\textsf{vec}(x_{i}x_{i}^{\top})_{j}$, where $\otimes$ is the Kronecker product of two vectors.
To protect the privacy of individual data elements, we rely on the notion of differential privacy, considering neighboring datasets as those differing by a single data point; see e.g.Dwork & Roth (2014) for an introduction. Recently, a series of works have developed effective methods for privately estimating individual matrices in the aforementioned product form by means of the the matrix mechanism (Li et al., 2015; Choquette-Choo et al., 2024; Denisov et al., 2022; Choquette-Choo et al., 2023b; Dvijotham et al., 2024; Kalinin & Lampert, 2024; Henzinger et al., 2024; Henzinger & Upadhyay, 2025; Fichtenberger et al., 2023; Henzinger et al., 2023; Choquette-Choo et al., 2023a; Kairouz et al., 2021; McMahan et al., 2024). Its core insight is that, in the continual release setting, outputs at later steps should require less noise to be made private than earlier ones because more data points contributed to their computation. One can exploit this fact by adding noise that is correlated between the steps instead of being independent. This way, at each step, some of the noise added at an earlier stage can be removed again, thereby resulting in a less noisy result without lowering the privacy guarantees.
Technically, the matrix mechanism relies on an invertible noise shaping matrix $C$. For our theoretical results, we often also assume that $C$ has decreasing column norms, As the private estimate of a product $AX$, one computes $\widehat{AX}=A(X+C^{-1}Z)$, where $Z$ is a matrix of i.i.d. Gaussian noise. This estimate is $(\epsilon,\delta)$-differentially private if the noise magnitude is at least $\text{sens}(CX)\cdot\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ denotes the variance required in the Gaussian distribution to ensure $(\epsilon,\delta)$-DP for sensitivity $1$ queries [^2] Here and in the following we do not provide a formula for $\sigma_{\epsilon,\delta}$, because closed-form expressions have been shown to be suboptimal in some regimes. Instead, we recommend determining its value numerically, such as in (Balle & Wang, 2018). For reference, typical values of $\sigma_{\epsilon,\delta}$ lie between approximately $0.5$ (low privacy, e.g. $\epsilon=8,\delta=10^{-3})$ and $50$ (high privacy, e.g. $\epsilon=0.1,\delta=10^{-9})$.. $\text{sens}(\cdot)$ denotes the sensitivity, which, for any function $F$, is defined as
$\text{sens}(F)=\max_{X\sim X^{\prime}}\|F(X)\!-\!F(X^{\prime})\|_{\text{F}}$ (2)
where $X\sim X^{\prime}$ denote any two data matrices, that are neighboring, i.e. identical except for one of the data vectors.
3 Joint Moment Estimation (JME)
Algorithm 1 Joint Moment Estimation (JME)
0:  stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$
0:  workload matrices $A_{1},A_{2}\in\mathbb{R}^{n\times n}$
0:  privacy parameters $(\epsilon,\delta)$
0:  noise shaping matrices $C_{1},C_{2}\in\mathbb{R}^{n\times n}$ (invertible with decreasing column norm) default: $\text{I}_{n\times n},\text{I}_{n\times n}$
  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-DP Gaussian mechanism
  $\lambda\leftarrow\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$ // scaling parameter,   where $c_{1}=\frac{8}{11+5\sqrt{5}}$, and $c_{d}=2$ for $d\geq 2$
  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // joint sensitivity
  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d^{2}}$ // 2nd moment noise
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
     yield   $\widehat{Y_{t}}=\sum\limits_{i=1}^{t}[A_{1}]_{t,i}\widehat{x_{i}},\ \ \widehat{S_{t}}=\sum\limits_{i=1}^{t}[A_{2}]_{t,i}\widehat{x_{i}\otimes x_{i}}$
  end for
We now introduce our main algorithmic contribution: the Joint Moment Estimation (JME) algorithm for solving the problem of differentially private (weighted) moment estimation in a continual release setting. Algorithm 1 shows its pseudo-code. JME takes as input a stream of input data vectors, the weights for the desired estimate in the form of two lower triangular workload matrices, and privacy parameters $\epsilon,\delta>0$. Optionally, it takes two noise-shaping matrices as input. If these are not provided, $C_{1}=C_{2}=\text{I}_{n\times n}$, can serve as defaults.
The algorithm uses the matrix mechanism to privately estimate the weighted sums of vectors for both the first and second moments. At each step, the algorithm receives a new data point, $x_{t}$, it creates private version of both $x_{t}$ and $x_{t}\otimes x_{t}$ by adding suitably scaled Gaussian noise. The noise shaping matrices, if provided, determine the covariance structure of the noise.
To balance between the estimates of the first moment and of the second moment, a scaling parameter, $\lambda$, is used. In Algorithm 1, this parameter is fixed such that the total sensitivity of estimating both moments is equal to the sensitivity of just estimating the first moment alone. This implies that the overall noise variance to make both moment estimates private is equal to that needed to achieve the same level of privacy for the first moment. Consequently, the fact that we also privately estimate the second moment does not increase the necessary noise level for the first moment, a property we call (second moment) privacy for free.
Note that privacy for free is a quite remarkable property of JME. Generally, when using the same data more than once for private computations, more noise for each of them is required to ensure the overall privacy, following, e.g., the composition theorems of DP (Kairouz et al., 2015). In some situations, however, one might also prefer a more flexible way to trade-off between the two moment estimates. For this, we present $\lambda$-JME in the appendix (Algorithm 3), a variant of JME that has $\lambda$ as a free hyper-parameter.
3.1 Properties of JME
For some inputs $X=(x_{1},\dots,x_{n})\in\mathbb{R}^{n\times d}$, let $Y=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $S=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the matrices of their first and second moments after each step, and let $\widehat{Y}=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $\widehat{S}=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the private estimates computed by Algorithm 1 with workload matrix $A$ and noise shaping matrix $C$. Then, the following properties hold.
Theorem 3.1 (Noise properties of JME).
$\widehat{Y}$ and $\widehat{S}$ are unbiased estimates of $Y$ and $S$. Their estimation noise, $\widehat{Y}-Y$ and $\widehat{S}-S$, has a symmetric distribution.
Theorem 3.2 (Privacy of JME).
Algorithm 1 is $(\epsilon,\delta)$-differentially private.
Theorem 3.3 (Utility of JME).
For any input $X$, the expected approximation error of Algorithm 1 for the first and second moments are
$\sqrt{\mathbb{E}\|Y-\widehat{Y}\|^{2}_{\text{F}}}$ $=2\zeta\sqrt{d}\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}\|A_{1}C_{1}^{-1}\|_{\text{F}}.$ (3)
$\sqrt{\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}}$ $=2\zeta\sqrt{c_{d}}d\sigma_{\epsilon,\delta}\|C_{2}\|_{1\to 2}\|A_{2}C_{2}^{-1}\|_{\text{F}},$ (4)
with $c_{d}$ as defined in Algorithm 1.
3.2 Proofs
Proof of Theorem 3.1.
The statement follows from the fact that for any $t=1,\dots,n$, JME’s estimates $\widehat{x_{t}}$ and $\widehat{x_{t}\otimes x_{t}}$ of $x_{t}$ and $x_{t}\otimes x_{t}$ are unbiased with symmetric noise distribution because they are constructed by adding zero-mean Gaussian noise. ∎
Sketch of the Proof of Theorem 3.2.
By the properties of the Gaussian mechanism, it suffices to show that the overall sensitivity of jointly estimating both moments has at most the value $s=2\zeta\|C_{1}\|_{1\to 2}$, as stated in Algorithm 1.
Definition 3.4.
For any noise-shaping matrices $C_{1}$ and $C_{2}$ and any $\lambda>0$, we define the joint sensitivity of estimating the first and $\lambda$-weighted second moment by
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})=\sup_{X\sim X^{\prime}}\!\Big{\|}\begin{pmatrix}C_{1}&\textbf{0}\\ \textbf{0}&\sqrt{\lambda}C_{2}\\ \end{pmatrix}\begin{pmatrix}X\!-\!X^{\prime}\\ X\operatorname{\!\bullet\!}X\!-\!X^{\prime}\operatorname{\!\bullet\!}X^{\prime}\end{pmatrix}\Big{\|}_{\text{F}}^{2}$ (5)
$=\sup_{X\sim X^{\prime}}\!\!\Big{[}\|C_{1}(X\!-\!X^{\prime})\|_{\text{F}}^{2}+\lambda\|C_{2}(X\operatorname{\!\bullet\!}X\!-\!X^{\prime}\operatorname{\!\bullet\!}X^{\prime})\|_{\text{F}}^{2}\Big{]}.$
The following Lemma (proved in the appendix) characterizes the values of the joint sensitivity as a function of $\lambda$.
Lemma 3.5 (Joint Sensitivity).
Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns. Then, for any $\lambda>0$ holds:
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}\left(\frac{\lambda\zeta^{2}\|C_{2}\|_{1\to 2}^{2}}{\|C_{1}\|_{1\to 2}^{2}}\right),$ (6)
where $\|\cdot\|_{1\to 2}$ denotes the maximum column norm, corresponding to the norm of the first column of $C_{1}$ and $C_{2}$, respectively. The function $r_{d}(\nu)$ is given by
$r_{d}(\nu)=\begin{cases}\frac{1}{8}(3-\tau)^{2}(\nu\tau+1+\nu),&\text{if }\nu>\frac{11+5\sqrt{5}}{8},\,d=1,\\[5.0pt] 2+2\nu+\frac{1}{2\nu},&\text{if }\nu>\frac{1}{2},\,d>1,\\[5.0pt] 4,&\text{otherwise,}\end{cases}$ (7)
where $\tau=\sqrt{1-\frac{2}{\nu}}$.
Algorithm 1 uses the scaling parameter $\lambda=\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$, so, by Lemma 3.5, the sensitivity of estimating both moment has the value $\sqrt{\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}(c_{d})}$. The choice of $c_{d}$ implies that $r_{d}(c_{d})=4$, which establishes that $\text{sens}_{\lambda}(C_{1},C_{2})=s$, which concludes the proof of Theorem 3.2. ∎
Sketch of the Proof of Theorem 3.3.
The identities follow from the general properties of the matrix mechanism. For any input $X$, the output of Algorithm 1 for the first moment is $\widehat{Y}=Y+A_{1}C^{-1}_{1}Z_{1}$, where $Z_{1}\sim[\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})]^{n\times d}$. Consequently, $\widehat{Y}-Y=A_{1}C^{-1}_{1}Z_{1}$, and hence
$\operatorname{\mathbb{E}}_{Z_{1}}\|\widehat{Y}-Y\|^{2}_{\text{F}}$ $=\|A_{1}C^{-1}_{1}\|^{2}_{\text{F}}\cdot\sigma^{2}_{\epsilon,\delta}s^{2}d$
Equation (3) follows by inserting $s=2\zeta\|C_{1}\|_{1\to 2}$. For the second moment, Equation (4) follows analogously using that the output of Algorithm 1 is $\widehat{S}=S+\lambda^{-\frac{1}{2}}A_{2}C^{-1}_{2}Z_{2}$ with $Z_{2}\sim[\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})]^{n\times d^{2}}$. ∎
3.3 Comparison with Alternative Techniques
Besides JME, a number of alternatives methods are possible that could be used to privately estimate the first and second moments. In this section, we introduce some of them and discuss their relation to JME.
3.3.1 Independent Moment Estimation (IME)
A straight-forward way to privately estimate first and second moments is to estimate and privatize both of them separately, where the necessary amount of noise is determined by the composition theorem of the Gaussian mechanism (Abadi et al., 2016) (see Algorithm 4 in the appendix).
IME resembles JME in the sense that (i) separate estimates of the moments are created, and (ii) the privatized results are unbiased estimators. However, it does not have JME’s privacy for free property. This is because IME relies on the composition theorem, so the privacy budget is split into two parts, one per estimate, where the exact split is a hyperparameter of the method. As a consequence, IME’s estimate of the first moment is always more noisy, and thereby of lower expected accuracy, than JME’s. For the second moment, IME could in principle achieve a lower noise than plain JME by adjusting the budget split parameter in an uneven way. This, however, would come at the expense of further increase in the error for estimating the first moment.
The following theorem establishes that $\lambda$-JME, the variant of JME with adjustable $\lambda$ parameter offers a strictly better trade-off than IME.
Theorem 3.6 (JME vs IME).
For any $\epsilon,\delta>0$, $\lambda$-JME Pareto-dominates IME with respect to the approximation error for the first vs second moment estimates.
The proof can be found in the appendix. Figure 1 visualizes this property graphically.
3.3.2 Concatenate-and-Split (CS)
In the special case where the two noise shaping matrices are meant to be the same (e.g. the trivial case where both are the identity matrix), it is possible to use a single privatization step for both moments. For this, one forms a new observation vector, $\tilde{x}_{i}$ by concatenating $x_{i}$ with a vectorized (and potentially rescaled version of) $x_{i}\otimes x_{i}$. Then, one privatizes the resulting vector, taking into account that $\tilde{x}_{i}\in\mathbb{R}^{d(d+1)}$ has higher dimension and a larger norm than the original $x_{i}\in\mathbb{R}^{d}$. The result is split again into first and second order components, and the latter is unscaled. The result are private estimate of $x_{t}$ and $x_{t}\otimes x_{t}$, from which the two weighted moment estimates can be constructed. Algorithm 5 in the appendix provides pseudocode for this concatenate-and-split (CS) method.
When applicable, CS is easy to implement and produces unbiased moment estimates. However, like IME, it has an unfavorable privacy-accuracy trade-off curve compared to JME, as formalized in the following theorem.
Figure 1: Approximation errors for the first and second moments ($d=10$, $n=100$, $C_{1}=C_{2}=I,A=E_{1}$, $\zeta=1$, and $\sigma=1$) across three methods: $\lambda$-JME with $\lambda$, IME with varying $\alpha$ parameters, and CS with varying $\tau$ parameter. One can see that $\lambda$-JME Pareto-dominates the other two methods.
Theorem 3.7 (JME vs CS).
For any $\epsilon,\delta>0$, $\lambda$-JME Pareto-dominates CS with respect to the approximation error for the first vs second moment estimates.
The proof can be found in the appendix. Figure 1 visualizes these cases (Theorems 3.6 and 3.7) graphically.
3.3.3 Post-processing (PP)
Post-processing (PP) is another easy-to-use method for joint moment estimation. It has appeared in the literature (Sheffet, 2019), at least in its naive form without the matrix factorization mechanism. For any $x_{i}$, it first computes a private estimate $\widehat{x_{t}}$ by adding sufficient noise to it. It then sets $\widehat{x_{t}\otimes x_{t}}:=\widehat{x_{t}}\otimes\widehat{x_{t}}$, which is automatically private by the postprocessing property of DP, and uses it to estimate the second moment matrix without additional privacy protection. Algorithm 6 in the appendix provides pseudocode.
(a) $d=1$, $C_{1}=C_{2}=\text{I}$
(b) $d=1$, $C_{1}=C_{2}=A^{1/2}$
(c) $d=1000$, $C_{1}=C_{2}=\text{I}$
(d) $d=1000$, $C_{1}=C_{2}=A^{1/2}$
Figure 2: Expected error of second moment estimation with JME versus PP with and without debiasing ($A=E_{1}$ (prefix sum), $n=1000$) In line with our analysis, for $d=1$ JME consistently achieves a higher quality than PP. For $d=1000$, JME is preferable to PP in the high privacy regime. Furthermore, the square root matrix factorization substantially improves the quality of both methods.
Like JME, PP has the privacy-for-free property, i.e., the private estimate of the second moment does not reduce the quality of the first moment. In contrast to JME, however, PP’s estimate of the second moments is not unbiased because the noise that was added to the first moment is squared during the process. It is possible to compensate for this by explicitly subtracting the bias, which can be computed analytically. The bias depends only on hyperparameters such as $n$, $d$, and $\sigma$, and not on the private data $X$ or the sampled noise used to protect it. Its exact expression can be found in equation (54).
For given privacy parameters, PP and JME use the same noise strength to privatize the first moment and therefore their estimates are of equal quality. For the second moment, PP and JME differ in their characteristics between the low privacy (small noise variance) and the high privacy (large noise variance) regime. To allow for a quantitative comparison, we derive characterizations of the approximation quality of PP.
Lemma 3.8 (Expected Second Moment Error for PP).
Assume the same setting as for Theorem 3.3, except that we compute the estimates $\widehat{Y}$ and $\widehat{S}_{\text{PP}}$ with the PP method. Then, the expected approximation error of the second moment satisfies:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}=\overbrace{d\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}^{\text{bias}}$
$\ \ +d(d+1)\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (8)
$\ \ +2(d+1)\sigma^{2}\zeta^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{\top}A_{2}\circ Q)XX^{T}),$
where $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}(Q)^{\top}$. For debiased PP, the term marked ”bias” does not occur.
In order to get a better impression of the relation between PP and JME, we first study the special case of a trivial factorization.
Corollary 3.9.
Assume that JME and PP use trivial factorizations, i.e. $C_{1}=C_{2}=I$. Then, JME’s expected approximation error for the second moment is
$\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}$ $=4c_{d}d^{2}\sigma^{2}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$ (9)
and the corresponding value for PP is
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=\big{(}d(d+1)\sigma^{4}+2(d+1)\sigma^{2}\big{)}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$
$\quad+\underbrace{d\sigma^{4}\zeta^{2}\|\sum_{i}(A_{2})_{[i,\cdot]}\|_{2}^{2}}_{\text{bias}},$ (10)
where $\sum_{i}(A_{2})_{[i,\cdot]}$ is the row-wise summation of $A_{2}$.
For debiased PP, the term marked ”bias” does not occur.
Proof. The proofs follow directly from Theorem 3.3 and Lemma 3.8 by observing that $\|C_{1}\|_{1\to 2}=\|C_{2}\|_{1\to 2}=1$.
In dimension $d=1$, JME has an error of $c_{1}\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ versus $(2\sigma^{4}+4\sigma^{2})\|A_{2}\|_{\text{F}}^{2}$ for (debiased) PP. Since $c_{1}<1$, for one-dimensional data, JME’s error is always lower than PP’s (see Figure 2 for visual confirmation). In high dimensions ($d\gg 1$), the terms quadratic in $d$ dominate, so the comparison is between $2\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ for JME and $d^{2}\sigma^{4}\|A_{2}\|_{\text{F}}^{2}$ for PP. Consequently, at least for $\sigma\geq\sqrt{2}$, JME achieves privacy with less added noise than PP.
In the regime of low privacy ($\sigma\to 0$) in high dimension $(d\gg 1)$, PP can be expected to result in higher accuracy estimates than JME, because the terms involving $\sigma^{4}$ make only minor contributions.
For settings with general noise shaping matrix, we cannot provide an exact comparison between PP and JME, because the $\sup_{X}$-term in Equation 8 has no closed-form solution. Instead, we derive upper and lower bounds (Lemma C.1 in the Appendix), and provide a numeric comparison in Figure 2.
unbiased? symmetric noise? privacy-for-free?
Method 1st / 2nd moment 1st / 2nd moment 2nd moment recommended regime of use
JME (proposed) ✓/ ✓ ✓/ ✓ high privacy or low dimensionality
IME ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
CS ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
PP ✓/✗ ✓/ ✗ low privacy in high dimensions
PP (Debiased) ✓/✓ ✓/ ✗ low privacy in high dimensions
Table 2: Overview of the properties for the methods discussed in Section 3. Properties unbiased and symmetric noise apply to each entry, stated separately for first and second moment. privacy-for-free applies only for the second moment.
4 Applications
To demonstrate potential uses of JME, we highlight two applications: private Gaussian density estimation, a classical probabilistic technique, and private Adam optimization, which is common in deep learning. The proofs for all theoretical results can be found in the appendix.
4.1 Private Gaussian density estimation
The maximum likelihood solutions to Gaussian density estimation from i.i.d. data, $x_{1},\dots,x_{t}$ famously is $\widehat{p}(x)=\mathcal{N}(x;\;\mu_{t},\Sigma_{t})$, where $\mu_{t}=\frac{1}{t}\sum_{i=1}^{t}x_{i}$ is the sample mean and $\Sigma_{t}=\frac{1}{t}\sum_{i=1}^{t}(x_{i}-\mu_{t})(x_{i}-\mu_{t})^{\top}$ is the sample covariance. Using the alternative identity $\widehat{\Sigma}_{t}=\frac{1}{t}\sum_{i=1}^{t}x_{i}x^{\top}_{i}-\mu_{t}\mu^{\top}_{t}$, one sees that the task can indeed be solved in a continuous release setting and that only estimates of the first and second moments are required.
(a) $d=1$, $n=1000$
(b) $d=100$, $n=1000$
Figure 3: Expected error of the estimated covariance matrix with JME versus PP (with trivial factorizations). For $d=1$, JME consistently achieves quality better than or on par with PP. For $d=100$, JME is preferable to PP only in the high privacy regime.
To do so privately, we use the averaging workload matrix $V=(a^{t}_{i})$ with $a^{t}_{i}=\frac{1}{t}$ for $1\leq i\leq t$ and $a^{t}_{i}=0$ otherwise, and we compute private estimates $\widehat{\mu}:=\widehat{VX}$ (private mean) and $\widehat{\Sigma}=\widehat{V(X\operatorname{\!\bullet\!}X)}-\widehat{(VX)}\operatorname{\!\bullet\!}\widehat{(VX)}$ (private covariance) using JME.
Note that $\widehat{\mu}$ is simply the first-moment vector as above. It is therefore unbiased and the guarantees of Theorem 3.3 holds for it. However, $\widehat{\Sigma}$ is not an unbiased estimate because in its computation the noise within $\widehat{VX}$ is squared. However, it is possible to characterize the bias analytically and subtract it if required. We provide the exact expression for the bias in equation (62) in the Appendix, where the debiased version is referred to as JME (Debiased).
The expected approximation error of $\widehat{\mu}$ is identical to the one in Theorem 3.3 with $A=V$ and $C_{1}=C_{2}=\text{I}$. The following Theorem establishes the approximation quality of the covariance estimate.
Theorem 4.1 (Private covariance matrix estimation with JME).
Assume that all input vectors have norm at most $1$. Let $\widehat{\Sigma}$ be the results of the above construction, where privacy is obtained by running JME with noise strength $\sigma$ and debiasing. Then it holds:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=(c_{d}d^{2}+2d+2)\sigma^{2}H_{n,1}$ (11)
$\quad+d(d+1)\sigma^{4}H_{n,2},$
with $c_{d}$ as in Theorem 3.3, and $H_{n,m}:=\sum\limits_{k=1}^{n}\frac{1}{k^{m}}$.
For comparison, we also analyze the case where the PP method is used to privatize the covariance matrix, $\widehat{\Sigma}_{\text{PP}}:=V(\widehat{X}\operatorname{\!\bullet\!}\widehat{X})-(V\widehat{X}\operatorname{\!\bullet\!}V\widehat{X})$, where $\widehat{X}$ are privatized entries of $X$. Again, the resulting biased estimate can be explicitly debiased, see Appendix (70) for the expression.
The following theorem states upper and lower bounds on the expected approximation error:
Theorem 4.2 (Private covariance matrix estimation with PP).
Assume the same setting as for Theorem 4.1. Let $\widehat{\Sigma}_{\text{PP}}$ be the result of the above construction, where privacy is obtained by running PP with noise strength $\sigma$ and debiasing. Then, for the expected error of the covariance matrix estimate it holds:
$d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}$
$\qquad+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
$\leq\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ (12)
$\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}.$
Comparison.
In the high privacy regime ($\sigma\gg 1$), the leading term for JME is $d(d+1)H_{n,2}\sigma^{4}$, and for PP it is $d(d+1)H_{n,1}\sigma^{4}$. Given that $H_{n,1}=O(\log n)$, while $H_{n,m}<\pi^{2}/6$ for $m\geq 2$, the error introduced by PP is logarithmically worse than JME. Figure 3 shows a numerical plot that confirms this observation. Note that our results match the lower bounds established by Kamath et al. (2020), who proved that private covariance estimation in the Frobenius norm requires $\Omega(d^{2})$ samples.
4.2 Private Adam optimization
The Adam optimizer (Diederik & Ba, 2015), has become a de-facto standard for optimization in deep learning. The defining property of Adam is its update rule, $\theta_{i}\leftarrow\theta_{i-1}-\alpha m_{i}/(\sqrt{v_{i}}+\epsilon)$, where $\alpha$ is a learning rate, and $m_{i}$ and $v_{i}$ are exponentially running averages of computed model gradients and componentwise squared model gradients, respectively. In the context of our work, these quantities correspond to a weighted first-moment vector and the diagonal of the weighted second-moment matrix.
Previous attempts to make Adam differentially private relied on postprocessing (Anil et al., 2022; Li et al., 2022), potentially with debiasing (Tang et al., 2024), i.e. they privatized the model gradients and derived the squared values from these. We demonstrate that JME’s approach of privatizing both quantities separately can be a competitive alternative. Algorithm 2 in the Appendix shows the pseudocode.
It contains some modifications compared to the original JME. In particular, we adjust JME to only estimate and privatize the diagonal of the second moment matrix, which reduces the runtime and memory requirements. Interestingly, as the next theorem shows, having to estimate only the diagonal elements of the covariance matrix does not reduce the problem’s sensitivity, so privatizing the estimates remains equally hard. This implies that our previous analyses, including the relation to the baselines, remain valid.
Theorem 4.3.
The sensitivity of JME, with the whole second-moment matrix $(C_{1}X,\sqrt{\lambda}C_{2}X\operatorname{\!\bullet\!}X)$, and with just the diagonal terms $(C_{1}X,\sqrt{\lambda}C_{2}X\!\circ\!X)$, are identical.
The proof can be found in the Section A.3. It does not follow from Lemma 3.5 and is significantly more intricate.
As in the previous cases, JME’s for free property ensures that its expected approximation error of the first moment is identical to PP. The following theorem establishes the expected approximation errors of both methods for the computed second moments (i.e. the diagonal of the second moment matrix):
Lemma 4.4 (Comparison of JME and PP for DP-Adam).
Let $D=(v_{1},\dots,v_{n})$ be the matrix of second moment estimates of the Adam algorithm. Denote by $\widehat{D}=(\widehat{v}_{1},\dots,\widehat{v}_{n})$ the private estimate of $D$ computed by Algorithm 2 with trivial factorization and noise strength $\sigma$, and let $\widehat{D}_{\text{PP}}$ be the analog quantity computed by DP-Adam with postprocessing. Then it holds:
$\mathbb{E}_{Z}\|\widehat{D}-D\|^{2}$ $=2d\sigma^{2}\cdot\|A_{2}\|_{\text{F}}^{2},$ (13)
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=(2d\sigma^{4}+4\sigma^{2})\cdot\|A_{2}\|_{\text{F}}^{2}$
$\quad+\underbrace{d\sigma^{4}\cdot\big{\|}\sum_{i}(A_{2})_{i}\big{\|}_{2}^{2}}_{\text{bias}},$ (14)
where $A_{2}$ is the workload matrix obtained from the coefficients of Adam’s exponentially weighted averaging operations. The term marked ”bias” disappears if PP is debiased.
The proof of the first part follows directly from Theorem 3.3. The error for PP is computed separately in Lemma C.4 in the appendix.
This lemma shows that as long as $\sigma\geq 1$, the error introduced by JME is strictly lower than that of PP (i.e. classic DP-Adam (Li et al., 2022)) and even the debiased version of PP (i.e. DP-AdamBC (Tang et al., 2024)). Figure 4 illustrates the relation in a prototypical case.
(a) trivial factorization
(b) square root factorization
Figure 4: Expected error of the estimated covariance vector for Adam with JME versus PP ($d=10^{6}$; $n=1000$). JME provides a better estimate in the mid to high privacy regime.
5 Experiments
Our main contributions in this work are both algorithmic and theoretical. Specifically, JME is the general purpose technique for moment estimation, which is promising for some scenarios and less promising for others. However, it is also a practical algorithm that can be easily implemented and integrated into standard machine learning pipelines. To demonstrate this, we report on the experimental result of using Algorithm 1 and Algorithm 2 in two exemplary settings, reflecting the application scenarios described above.
Private Gaussian density estimation.
From a given data distribution, $p(x)=\mathcal{N}(\mu,\Sigma)$, we sample $n=200$ data points and use either JME or PP to form a private estimates, $\widehat{p}_{t}(x)=\mathcal{N}(\widehat{\mu}_{t},\widehat{\Sigma}_{t})$, at each step $t=1,\dots,n$, of the continuous release process. To ensure positive definiteness, we symmetrize JME’s estimated covariance matrices and project them onto the positive definite cone. As a postprocessing operation, this does not affect their privacy.
Figure 5 shows the results, with the approximation quality, measured by the Kullback-Leibler (KL) divergence, $\text{KL}(\widehat{p}_{t}\|p)$, at each step, $t$, as average and standard deviation over 1000 runs, with $\mu\sim\mathcal{N}(0,\frac{1}{2}I_{d})$ and $\Sigma\sim W_{d}\left(\frac{1}{2}I_{d},2d\right)$ (i.e. a Wishart distribution) in each case. One can see that in the high-privacy regime (here: $\sigma=2$), on average, JME achieves a better estimate of the true density than PP, with and without debiasing.
Figure 5: Approximation quality of JME and PP for private Gaussian density estimation ($d=5,n=100$) as average and standard deviation over 1000 runs. In the high-privacy regime ($\sigma=2$), JME achieves lower KL divergence than PP after $\approx 10$ samples.
(a) batchsize $B\!=\!256$, $\varepsilon\approx 1.7$
(b) batchsize $B\!=\!1$, $\varepsilon\approx 0.16$
Figure 6: Results of private Adam training on CIFAR-10 experiments comparing four methods: DP-Adam based PP (with and without debiasing), JME, and joint clipping, over 10 epochs. Left: for low to medium privacy, JME outperforms vanilla DP-Adam but is slightly worse than the debiased version. Right: for high privacy, JME is slightly better than the debiased version, where as vanilla DP-Adam cannot handle the necessarily amounts of noise.
Private model training with Adam.
We train a convolutional network on the CIFAR-10 dataset with DP-Adam, which is privatized either with JME or PP. Because gradients could have arbitrarily large number, we apply gradient clipping to a model-selected threshold in both methods. In addition, we include a heuristic baseline that uses the concatenate-and-split techniques in which not the norm of the gradient vector but the norm of the concatenation vector is clipped. While inferior to JME in the worst case, this might be beneficial for real data, so we include it in the experimental evalution.
The results in Figure 6 confirm our expectations: in a setting with small noise variance (large batchsize, low privacy), DP-Adam-JME achieved better results than standard DP-Adam, but worse than the debiased variant of DP-Adam. If the noise variance is large (small batchsize, high privacy), DP-Adam-JME slightly improves over the other methods. For detailed accuracy results, see Table 3 in the appendix with hyperparameters provided in Table 4.
6 Related Works
The problem of differentially private moments (and the related problem of covariance estimation) has a rich history of development, with optimal results known in the central model of privacy (Smith, 2011) as well as the local model of privacy (Duchi et al., 2013), and also in the worst-case setting. Sheffet (2019) proposes three differentially private algorithms for approximating the second-moment matrix, each ensuring positive definiteness. The related setting of covariance estimation has been studied in the worst-case setting when the data comes from a bounded $\ell_{2}$ ball by several works for approximate differential privacy (Achlioptas & McSherry, 2007; Blum et al., 2005; Dwork et al., 2014; Mangoubi & Vishnoi, 2022, 2023; Upadhyay, 2018) with the works of Amin et al. (2019) and Kapralov & Talwar (2013) presenting a pure differentially private algorithm.
Covariance estimation is also used as a subroutine in mean-estimation work under various distributional assumptions (Kamath, 2024). However, none of these approaches are directly applicable to the continual release model and they offer improvements over the Gaussian mechanism only in a very high privacy regime.
7 Summary and Discussion
In this work we studied the problem of jointly estimating the first and second moment of a continuous data stream in a differentially private way. We presented the Joint Moment Estimation (JME) method, which solves the problem by exploiting the recently proposed matrix mechanism with carefully tuned noise level. As a result, JME produces unbiased estimates of both moments while requiring less noise to be added than baseline methods, at least is the high privacy regime. We applied JME to private Gaussian density estimation and model training with Adam, demonstrating improved performance in high-privacy regimes both theoretically and practically.
Despite the promising results, several open question remain. In particular, would like to explore if postprocessing is indeed the optimal strategy in a low-privacy regime, or if a better privacy-utility trade-off is still possible. Furthermore, we plan to explore the possibility of problem-specific factorizations, which could be fused with the proposed method.
Acknowledgments
We thank Monika Henzinger for her valuable feedback and insightful discussions on earlier versions of this draft. We are also grateful to Mher Safaryan for his contributions to discussions on DP-Adam. Additionally, we thank Ryan McKenna for suggesting Joint Clipping (CS) as a baseline. Jalaj Upadhyay’s research was funded by the Rutgers Decanal Grant no. 302918 and an unrestricted gift from Google. A part of this work was done while visiting the Institute of Science and Technology Austria (ISTA). Christoph Lampert’s research was partially funded by the Austrian Science Fund (FWF) 10.55776/COE12.
Appendix A Proofs of the Main Theorems
In this section, we provide proofs of Theorem 3.2 and Theorem 3.3 from Section 3 as well as Theorem 4.3 from Section 4.
A.1 Proof of Theorem 3.2 and Lemma 3.5
The privacy of Algorithm 1 follows from the properties of matrix mechanism (Li et al., 2015), with a precise estimate of the sensitivity of the joint estimation process that we will introduce and discuss later in this section.
By means of the matrix factorization mechanism with $A_{1}=B_{1}C_{1}$ and $A_{2}=B_{2}C_{2}$, we write the joint moment estimate as,
$\begin{pmatrix}Y&\boldsymbol{0}\\ \boldsymbol{0}&S\end{pmatrix}$ $=\begin{pmatrix}B_{1}&\boldsymbol{0}\\ \boldsymbol{0}&\lambda^{-\frac{1}{2}}B_{2}\end{pmatrix}\begin{pmatrix}C_{1}X&\boldsymbol{0}\\ \boldsymbol{0}&\lambda^{\frac{1}{2}}C_{2}(X\operatorname{\!\bullet\!}X)\end{pmatrix}$ (15)
where $\lambda>0$ is an arbitrary trade-off parameter that we will adjust optimally later. To make the estimate private, we privatize the rightmost matrix in (15), which contains the data, by adding suitable scaled Gaussian noise. The subsequent matrix multiplication then acts on private data, so its result is also private.
It remains to show that the noise level specified in Algorithm 1 suffices to guarantee $(\epsilon,\delta)$-privacy.
For this, we denote by $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ be the set of input sequences where each $d$-dimensional vectors has bounded $\ell_{2}$ norm. The sensitivity of a computation is the maximal amount by which the result differs between two input sequences $X,X^{\prime}\in\mathcal{X}$, which are identical except for a single data vector at an arbitrary index (denoted by $X\sim X^{\prime}$).
For JME, the relevant sensitivity is the one of the matrix we want to privatize, i.e. (in the squared form)
$\operatorname{sens}_{\lambda}^{2}(C_{1},C_{2})\!=\!\!\sup_{X\sim X^{\prime}}\left\|\begin{pmatrix}C_{1}&\textbf{0}\\ \textbf{0}&\sqrt{\lambda}C_{2}\\ \end{pmatrix}\begin{pmatrix}X-X^{\prime}\\ X\operatorname{\!\bullet\!}X\!-\!X^{\prime}\operatorname{\!\bullet\!}X^{\prime}\end{pmatrix}\right\|_{\text{F}}^{2}$
$\quad=\sup_{X\sim X^{\prime}}\!\!\!\left[\|C_{1}(X\!-\!X^{\prime})\|_{\text{F}}^{2}+\lambda\|C_{2}(X\operatorname{\!\bullet\!}X\!-\!X^{\prime}\operatorname{\!\bullet\!}X^{\prime})\|_{\text{F}}^{2}\right]$ (16)
Due to the linearity of the operations and the condition imposed by $X\sim X^{\prime}$, most terms in (16) cancel out and the value of $\operatorname{sens}^{2}_{\lambda}(C_{1},C_{2})$ simplifies into the solution of the following optimization problem.
Problem 1 (Sensitivity for Joint Moment Estimation).
$\max_{i=1,\dots,n}\max_{\begin{subarray}{c}\|x\|_{2}\leq\zeta\\ \|y\|_{2}\leq\zeta\end{subarray}}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2},$ (17)
where $\alpha^{2}_{i}=\|(C_{1})_{i}\|_{2}^{2}$ and $\beta^{2}_{i}=\|(C_{2})_{i}\|_{2}^{2}$ are the column norms of the matrices $C_{1}$ and $C_{2}$, respectively.
To study Problem 1, we introduce as an intermediate object the formulation of (17) in the special case of $\zeta=1$ and $\alpha_{i}=\beta_{i}=1$ for $i=1,\dots,n$, treated as a function of $\lambda$:
$r_{d}(\lambda):=\max_{\begin{subarray}{c}x,y\in\mathbb{R}^{d}\\ \|x\|_{2}\leq 1\\ \|y\|_{2}\leq 1\end{subarray}}\|x-y\|^{2}_{2}+\lambda\|x\otimes x-y\otimes y\|_{\text{F}}^{2}.$ (18)
The following theorems provide specific values for $r_{d}$ in the special case of $d=1$ (Theorem A.1) and for general $d\geq 2$ (Theorem A.2):
Theorem A.1 ($d=1$).
For any $\lambda>0$, it holds:
$r_{1}(\lambda)\!=\!\begin{cases}4&\text{if }\lambda\leq\frac{11+5\sqrt{5}}{8},\\ \frac{1}{8}(3-\tau)^{2}(\lambda\tau+1+\lambda)&\text{otherwise.}\end{cases}$ (19)
where $\tau=\sqrt{1-2/\lambda}$. Moreover, the function $r_{1}(\lambda)$ is a continuous function with respect to the parameter $\lambda>0$.
Theorem A.2 (Joint Sensitivity for Moments Estimation).
For $d>1$ and $\lambda>0$:
$r_{d}(\lambda)$ =$=\begin{cases}4\quad&\text{if}\quad\lambda\leq\frac{1}{2},\\ 2+2\lambda+\frac{1}{2\lambda}\quad&\text{otherwise.}\\ \end{cases}$ (20)
The proofs of both theorems can be found further in the appendix. Theorem A.1 requires only straightforward optimization. For theorem A.2, we rewrite the Frobenius norm of the difference of Kronecker products and optimize over all possible values for $\langle x,y\rangle$.
As a corollary of Theorems A.1 and A.2, we obtain a characterization of the general solutions to Problem 1.
Corollary A.3 (Solution to Problem 1).
Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns. Then, for any scaling parameter $\lambda>0$, it holds that
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\zeta^{2}\alpha^{2}_{1}r_{d}(\lambda\zeta^{2}\beta^{2}_{1}/\alpha^{2}_{1})$ (21)
where $r_{d}$ is specified in (19) or (20), and $\alpha^{2}_{1}$ and $\beta^{2}_{1}$ are the squared norms of the first columns of the matrices $C_{1}$ and $C_{2}$, respectively.
Proof.
A straightforward calculation shows
$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\max_{i=1,\dots,n}\sup_{\|x\|\leq\zeta,\|y\|\leq\zeta}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2}$
$\;=\frac{\zeta^{2}}{\alpha^{2}_{1}}\left[\sup_{\|x\|\leq 1,\|y\|\leq 1}\|x-y\|_{2}^{2}+\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\|x\otimes x-y\otimes y\|_{2}^{2}\right]$
$\;=\zeta^{2}\alpha^{2}_{1}r_{d}(\lambda\zeta^{2}\beta^{2}_{1}/\alpha^{2}_{1})$
completing the proof of Corollary A.3. ∎
Corollary 1 implies that the joint estimate (15) will be $(\epsilon,\delta)$-private, if we use noise of strength at least, $\sigma=\zeta^{2}\alpha_{1}r_{d}(\lambda\zeta^{2}\beta_{1}/\alpha_{1})\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ is the noise strength required for the Gaussian mechanism with sensitivity $1$. This finishes the proof of the Lemma 3.5.
The claim of Theorem 3.2 follows, because Algorithm 1 corresponds exactly to the above construction, only making use of the identity $B(CX+Z)=A(X+C^{-1}Z)$, and for the special case of $\lambda:=\lambda^{*}$, defined as
$\lambda^{*}:=\frac{\alpha^{2}_{1}}{\beta^{2}_{1}\zeta^{2}c_{d}},$ (22)
with $c_{d}=2$ if $d>1$ and $\frac{8}{11+5\sqrt{5}}$ for $d=1$, which is the smallest values for $\lambda$, such that $r_{d}\left(\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\right)=4$. The sensitivity is then equal to $\text{sens}_{\lambda^{*}}(C_{1},C_{2})=2\zeta\alpha_{1}=2\zeta\|C_{1}\|_{1\to 2}$, where the last identity holds because of $C_{1}$’s decreasing column norm structure.
We furthermore note that this value is exactly the sensitivity of estimating the first moment alone, because
$\max_{X\sim X^{\prime}}\|C_{1}X\!-\!C_{1}X^{\prime}\|^{2}\!=\!\max_{i=1,\dots,n}\max_{\begin{subarray}{c}\|x\|\leq\zeta\\ \|y\|\leq\zeta\end{subarray}}\alpha_{i}\|x-y\|^{2}=4\alpha_{1}^{2}\zeta^{2}$ (23)
This means that JME estimates the second moment without increasing the noise for the first moment, proving our claim that we obtain the second moment privacy for free.
A.2 Proof of Theorem 3.3
To prove Theorem 3.3, we have to determine how the left-hand side of (15) changes in expectation due to the added noise on the right-hand side. Due to the additive nature of the moment estimation process, we can do so explicitly.
For $Z_{1}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d}$ it holds that
$\mathbb{E}_{Z}\|Y-\hat{Y}\|^{2}_{\text{F}}$ $=\mathbb{E}_{Z}\|B_{1}Z_{1}\|^{2}=d\sigma^{2}\|B_{1}\|^{2}_{\text{F}}.$ (24)
Inserting $B_{1}=A_{1}C_{1}^{-1}$ and $\sigma=2\zeta\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}$, Equation (3) follows. Analogoulsy, for $Z_{2}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d\times d}$,
$\mathbb{E}_{Z}\|S-\hat{S}\|^{2}_{\text{F}}$ $=\mathbb{E}_{Z}\|\frac{1}{\sqrt{\lambda^{*}}}B_{2}Z_{2}\|^{2}=\frac{d^{2}\sigma^{2}}{\lambda^{*}}\|B_{2}\|^{2}_{\text{F}}.$ (25)
With $\sigma$ as above, Equation (4) follows from $B_{2}=A_{2}C_{2}^{-1}$, and JME’s specific choice of $\lambda^{*}=\frac{\|C_{1}\|^{2}_{1\to 2}}{\|C_{2}\|^{2}_{1\to 2}\zeta^{2}c_{d}}$.
A.3 Proof of Theorem 4.3
The proof of Theorem 4.3 goes similarly to 3.2, but now we estimate only the diagonal of the second moment matrix. We will show that the sensitivity remains unchanged by proving that the corresponding function $r_{d}(\lambda)$ in (18) also remains unchanged. Specifically:
$r^{\text{diag}}_{d}(\lambda)=\sup\limits_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\!\!\!\big{[}\|x-y\|_{2}^{2}+\lambda\|\operatorname{\operatorname{diag}}(x\otimes x)-\operatorname{\operatorname{diag}}(y\otimes y)\|_{2}^{2}\big{]}=r_{d}(\lambda)$
For simplicity, we denote $x\circ x=\operatorname{\operatorname{diag}}(x\otimes x)$.
In dimension $d=1$, the two functions are identical by construction. In dimension $d=2$, we compute the new function explicitly using the following lemma:
Lemma A.4.
Consider $x,y\in\mathbb{R}^{2}$, and let $\lambda>0$ then,
$r_{2}^{\text{diag}}(\lambda)=\sup_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\left[\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\right]=\begin{cases}4,&\text{if }\lambda\leq\frac{1}{2},\\ 2+2\lambda+\frac{1}{2\lambda},&\text{if }\lambda>\frac{1}{2}.\end{cases}$ (26)
Next, we use a dimension reduction argument to prove that $r^{\text{diag}}_{d}(\lambda)=r^{\text{diag}}_{2}(\lambda)$ for all $d\geq 2$, via the following lemma:
Lemma A.5 (Dimension Reduction).
For any vectors $x,y\in\mathbb{R}^{d}$, where $d\geq 3$, there exist vectors $x^{\prime},y^{\prime}\in\mathbb{R}^{d-1}$ that for any $\lambda>0$ satisfies the inequality:
$\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\leq\|x^{\prime}-y^{\prime}\|_{2}^{2}+\lambda\|x^{\prime}\circ x^{\prime}-y^{\prime}\circ y^{\prime}\|_{2}^{2}.$
We apply this lemma recursively to prove that $r^{\text{diag}}_{d}(\lambda)=r^{\text{diag}}_{2}(\lambda)$ for all $d\geq 2$. The proof of the lemma can be found later in the appendix.
By combining these lemmas, we conclude the proof of the theorem.
Appendix B Additional Materials
(a) $d=100,\beta=0.9$, $C_{1}=C_{2}=I$
(b) $d=100,\beta=0.9$, $C_{1}=C_{2}=A_{\beta}^{1/2}$
(c) $d=100,w=10,$ $C_{1}=C_{2}=I$
(d) $d=100,w=10,$ $C_{1}=C_{2}=A_{w}^{1/2}$
Figure 7: Expected error of second moment estimation with JME versus PP under different workload settings. The scenarios include exponential decay ($\beta=0.9$) and sliding window ($w=10$) workloads and $d=100,n=1000$, both trivial and square root matrix factorization. In line with our analysis, incorporating matrix factorization significantly improves the quality of both methods, particularly in the high privacy regime.
Method Epoch 1 Epoch 2 Epoch 3 Epoch 4 Epoch 5 Epoch 6 Epoch 7 Epoch 8 Epoch 9 Epoch 10
$\varepsilon\approx 0.16$ DP-Adam-JME 12.43 $\pm$ 3.95 19.18 $\pm$ 2.27 23.29 $\pm$ 1.25 25.83 $\pm$ 0.48 27.25 $\pm$ 0.20 27.81 $\pm$ 0.14 28.50 $\pm$ 0.40 28.89 $\pm$ 0.54 29.61 $\pm$ 0.51 30.38 $\pm$ 0.62
DP-Adam-Clip 13.54 $\pm$ 3.31 19.68 $\pm$ 2.76 24.92 $\pm$ 1.48 26.24 $\pm$ 1.04 26.60 $\pm$ 0.25 27.27 $\pm$ 0.55 27.56 $\pm$ 0.46 27.98 $\pm$ 0.26 27.78 $\pm$ 0.17 28.14 $\pm$ 0.33
DP-Adam-Debiased 12.41 $\pm$ 4.14 16.91 $\pm$ 1.36 23.78 $\pm$ 0.70 24.92 $\pm$ 0.55 26.20 $\pm$ 0.65 26.96 $\pm$ 1.19 27.46 $\pm$ 0.90 28.14 $\pm$ 0.72 28.55 $\pm$ 0.92 29.14 $\pm$ 0.60
DP-Adam 10.29 $\pm$ 0.51 10.32 $\pm$ 0.56 10.32 $\pm$ 0.56 10.34 $\pm$ 0.58 10.39 $\pm$ 0.67 10.41 $\pm$ 0.70 10.42 $\pm$ 0.73 10.45 $\pm$ 0.79 10.45 $\pm$ 0.78 10.48 $\pm$ 0.83
$\varepsilon\approx 1.7$ DP-Adam-JME 32.64 $\pm$ 0.71 39.19 $\pm$ 1.52 42.28 $\pm$ 0.48 $4$$\pm$ 0.57 45.12 $\pm$ 0.81 45.61 $\pm$ 0.73 46.91 $\pm$ 1.21 47.04 $\pm$ 0.98 47.24 $\pm$ 0.79 47.94 $\pm$ 0.85
DP-Adam-Clip 30.51 $\pm$ 0.72 38.03 $\pm$ 1.00 40.66 $\pm$ 0.75 42.88 $\pm$ 0.46 43.30 $\pm$ 0.71 43.52 $\pm$ 0.50 44.51 $\pm$ 0.70 44.45 $\pm$ 0.24 44.76 $\pm$ 0.22 44.34 $\pm$ 0.28
DP-Adam-Debiased 34.21 $\pm$ 0.95 41.38 $\pm$ 1.54 45.20 $\pm$ 0.44 46.91 $\pm$ 1.25 48.06 $\pm$ 0.36 48.38 $\pm$ 0.83 48.26 $\pm$ 1.05 48.38 $\pm$ 1.21 47.81 $\pm$ 0.73 48.53 $\pm$ 0.99
DP-Adam 26.58 $\pm$ 0.43 31.45 $\pm$ 0.43 35.02 $\pm$ 1.16 36.39 $\pm$ 0.58 38.15 $\pm$ 0.95 39.31 $\pm$ 0.52 40.33 $\pm$ 0.33 41.47 $\pm$ 1.05 42.42 $\pm$ 0.94 43.07 $\pm$ 0.86
Table 3: CIFAR-10 experiments with two different privacy budgets, $\varepsilon\approx 1.7$ and $\varepsilon\approx 0.16$ for $\delta=10^{-6}$, for four methods: DP-Adam with and without debiasing, JME, and Joint Clipping. The average and standard deviation errors are based on 3 runs.
Noise Mult Batch Size Method lr Scaling ($\lambda,\tau)$ eps Clip Norm
$\varepsilon\approx 1.7$ 1 256 DP-Adam-JME $10^{-3}$ 1 $10^{-6}$ 1
DP-Adam-Clip $10^{-3}$ 0.5 $10^{-6}$ 1
DP-Adam-Debiased $10^{-3}$ $10^{-6}$ 1
DP-Adam $10^{-3}$ $10^{-8}$ 1
$\varepsilon\approx 0.16$ 2 1 DP-Adam-JME $10^{-7}$ 1 $10^{-7}$ 1
DP-Adam-Clip $10^{-7}$ 0.5 $10^{-7}$ 1
DP-Adam-Debiased $10^{-7}$ $10^{-7}$ 1
DP-Adam $10^{-7}$ $10^{-8}$ 1
Table 4: Hyperparameters for CIFAR-10 Experiments. Medium-privacy experiments use a batch size of $256$, compared to $1$ in the high-privacy regime, while using a noise multiplier of $\sigma_{\varepsilon,\delta}=1$ for the medium-privacy regime and $\sigma_{\varepsilon,\delta}=2$ for the high-privacy regime. JME and joint clipping require an additional hyperparameter—scaling—which is optimized to find the best value for those runs. We also find it helpful to clip the updates; for this, we use the same clipping norm.
Definition B.1 (Face-Splitting Product).
Let $A=(a_{i}^{\top})_{i=1}^{n}$ with $a_{i}\in\mathbb{R}^{d_{1}}$ and $B=(b_{i}^{\top})_{i=1}^{n}$ with $b_{i}\in\mathbb{R}^{d_{2}}$. The Face-Splitting Product of $A$ and $B$, denoted by $A\operatorname{\!\bullet\!}B$, is defined as:
$A\operatorname{\!\bullet\!}B=(a_{i}\otimes b_{i})_{i=1}^{n}=(\operatorname{vec}(a_{i}b_{i}^{\top}))_{i=1}^{n},$ (27)
where $\otimes$ denotes the Kronecker product, and $\operatorname{vec}$ denotes vectorization.
The result is a matrix of size $n\times(d_{1}d_{2})$, where each row corresponds to the vectorized Kronecker product of $a_{i}$ and $b_{i}$. A few properties of the Face-Splitting Product that we will use further:
  • Bilinearity $A\operatorname{\!\bullet\!}(B+C)=A\operatorname{\!\bullet\!}B+A\operatorname{\!\bullet\!}C$ and $(A+B)\operatorname{\!\bullet\!}C=A\operatorname{\!\bullet\!}C+B\operatorname{\!\bullet\!}C)$.
  • Associativity $(A\operatorname{\!\bullet\!}B)\operatorname{\!\bullet\!}C=A\operatorname{\!\bullet\!}(B\operatorname{\!\bullet\!}C)$.
  • Frobenious norm $\|A\operatorname{\!\bullet\!}B\|_{\text{F}}=\|B\operatorname{\!\bullet\!}A\|_{\text{F}}$, even so $A\operatorname{\!\bullet\!}B\neq B\operatorname{\!\bullet\!}A$.
Appendix C Technical Proofs
C.1 Bounds on the expected approximation error for PP with non-trivial factorization
Lemma C.1.
$\|A_{2}C_{1}^{-1}\|^{2}_{\text{F}}$ $\leq\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}\big{(}(A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top})XX^{\top}\big{)}\leq\sum_{ij}\big{|}\,[A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top}]_{i,j}\,\big{|}$ (28)
Proof of Lemma C.1.
The proof is elementary. For the lower bound. consider the specific choice for $X=(x_{1},\dots,x_{n})$ with all rows identical with $\|x_{i}\|=1$, such that $XX^{\top}$ is the constant matrix of all $1$s. For the upper bound, we observe that $[XX^{\top}]_{i,j}\in[-1,1]$, so
$\operatorname{\operatorname{tr}}\big{(}(A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top})XX^{\top}\big{)}$ $=\sum_{i,j}[A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top}]_{i,j}[XX^{\top}]_{i,j}\leq\sum_{i,j}\big{|}\,[A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top}]_{i,j}\,\big{|}.$ (29)
See A.1
Proof.
We recall that the function $r_{1}(\lambda)$ is defined as
$r_{1}(\lambda)=\sup_{|x|,|y|\leq 1}[(x-y)^{2}+\lambda(x^{2}-y^{2})^{2}].$ (30)
Given the difference it is an increasing function of $x+y$ therefore without any loss of generality we can assume $y=1$. Then the derivative of this expression would give us
$2(x-1)(1+2\lambda(x^{2}+x))$ (31)
The optimal value will depend on the roots of the quadratic polynomial. The determinant is $4\lambda^{2}-8\lambda$ therefore for $\lambda<2$ there are no roots, the function is decreasing on the segment $[-1,1]$ reaching its maximal value at $x=-1$ with the value $4$. Otherwise we have the following roots $\frac{-1\pm\sqrt{1-2/\lambda}}{2}$. We observe that for large $\lambda\gg 1$ we have the optimal $x=0$ which corresponds to the maximal value for the second term. The maximum could be in both $x=-1$ and $x=\frac{-1+\sqrt{1-2/\lambda}}{2}$ therefore we will need to compare them. By substituting it into the expression we get $\frac{1}{8}(\tau-3)^{2}(\lambda\tau+1+\lambda)$, where $\tau=\sqrt{1-2/\lambda}$. We should compare this expression with $4$ to find a boundary when $x=-1$ is an optimal solution.
$\frac{1}{8}(\tau-3)^{2}(\lambda\tau+1+\lambda)\leq 4$ (32)
We can express $1/\lambda=\frac{1-\tau^{2}}{2}$. Therefore we get the following inequality of variable $10\leq\tau<$ 1:
$\begin{split}\frac{1}{8}(\tau-3)^{2}(\tau+\frac{1-\tau^{2}}{2}+1)-4\frac{1-\tau^{2}}{2}&=\frac{1}{16}(3-\tau)^{3}(1+\tau)-2(1-\tau)(1+\tau)\\ \quad&=\frac{1}{16}(1+\tau)\left[27-27\tau+9\tau^{2}-\tau^{3}-32+32\tau\right]\\ \quad&=-\frac{1}{16}(1+\tau)^{2}(\tau^{2}-10\tau+5)\leq 0\end{split}$ (33)
Which is less than 00 for $\tau\leq 5-2\sqrt{5}\Rightarrow\lambda\leq\frac{11+5\sqrt{5}}{8}\approx 2.77$ which concludes the proof. ∎
See A.2
Proof.
First, we decompose the Kronecker product as follows:
$\|x\otimes x-y\otimes y\|_{\text{F}}^{2}$ $=\|xx^{T}-yy^{T}\|_{\text{F}}^{2}=\operatorname{\operatorname{tr}}(xx^{T}xx^{T})-2\operatorname{\operatorname{tr}}(xx^{T}yy^{T})+\operatorname{\operatorname{tr}}(yy^{T}yy^{T})$ (34)
$=\|x\|_{2}^{4}+\|y\|_{2}^{4}-2\langle x,y\rangle^{2}.$
Similarly, the squared norm of the difference is
$\|x-y\|_{2}^{2}=\|x\|_{2}^{2}+\|y\|_{2}^{2}-2\langle x,y\rangle.$ (35)
Therefore, the objective function becomes
$\|x-y\|^{2}+\lambda\|x\otimes x-y\otimes y\|_{\text{F}}^{2}=\|x\|_{2}^{2}+\|y\|_{2}^{2}+\lambda\|x\|_{2}^{4}+\lambda\|y\|_{2}^{4}-2\langle x,y\rangle-2\lambda\langle x,y\rangle^{2}.$ (36)
The first four terms are maximized when $\|x\|=\|y\|=1$, yielding $2+2\lambda$. The last two terms can then be optimized independently over the scalar product $\langle x,y\rangle$ in dimension $d>1$.
Let $\beta=\langle x,y\rangle$, where $-1\leq\beta\leq 1$. The expression $-\beta-2\lambda\beta^{2}$ is maximized at $\beta=-\frac{1}{2\lambda}$ when $\lambda\geq\frac{1}{2}$, or at $\beta=-1$ when $\lambda<\frac{1}{2}$. If $\beta=-1$, then $x=-y$ and the objective function becomes $4$. If $\beta=-\frac{1}{2\lambda}$, the objective function is equal to $2+2\lambda+\frac{1}{2\lambda}$, which concludes the proof.
See 3.6
Proof.
We begin the proof with the following observation: $\lambda$-JME and IME introduce additive noise to the first and second moments. Therefore, for the Frobenius approximation error, it is sufficient to compare just the variances of the noise introduced by those methods. We use an instance of a Gaussian mechanism, and the privacy guarantees are more appropriately characterized by the notion of Gaussian privacy. For the sake of the proof, we assume that $(\epsilon,\delta)$-DP is equivalent to $\mu$-GDP for a specific choice of $\mu$. The sensitivity of $\lambda$-JME depends on the dimensionality. Here, we assume $d>1$ to address the hardest case, as $d=1$ follows trivially.
To estimate $x$ and $x\otimes x$ simultaneously in a differentially private way via composition theorem, we need to split the privacy budget between the components. Using the Gaussian mechanism, we split the budget as $\mu_{1}$-GDP for the first component and $\mu_{2}$-GDP for the second component such that $\mu_{1}^{2}+\mu_{2}^{2}=\mu^{2}$. The squared sensitivity of $x$ is $4\zeta^{2}$, and the squared sensitivity of $x\otimes x$ is $2\zeta^{4}$, assuming $\|x\|_{2}\leq\zeta$. Therefore, the variance of noise added to the first and second components is given by:
$\left(\frac{4\zeta^{2}}{\mu_{1}^{2}},\frac{2\zeta^{4}}{\mu^{2}-\mu_{1}^{2}}\right).$ (37)
Our analysis for JME yields the pair of variances, for $\lambda\zeta^{2}\geq\frac{1}{2}$:
$\left(\frac{\zeta^{2}(2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}})}{\mu^{2}},\frac{\zeta^{2}(2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}})}{\lambda\mu^{2}}\right).$ (38)
We aim to show that for a given variance in the first component, our variance for the second component is smaller. Specifically, we need to prove:
$\frac{\zeta^{2}(2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}})}{\lambda\mu^{2}}<\frac{2\zeta^{4}}{\mu^{2}-\mu_{1}^{2}},\quad\text{where}\quad\frac{4\zeta^{2}}{\mu_{1}^{2}}=\frac{\zeta^{2}(2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}})}{\mu^{2}}.$ (39)
By substituting $\frac{\mu^{2}}{\mu_{1}^{2}}=\frac{1}{2}+\frac{\lambda\zeta^{2}}{2}+\frac{1}{8\lambda\zeta^{2}}$ into the inequality, we obtain:
$\frac{2\lambda\zeta^{2}}{2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}}}>1-\frac{1}{\frac{1}{2}+\frac{\lambda\zeta^{2}}{2}+\frac{1}{8\lambda\zeta^{2}}}.$ (40)
Multiplying both sides by the denominator yields:
$42\lambda\zeta^{2}>2+2\lambda\zeta^{2}+\frac{1}{2\lambda\zeta^{2}}-$ 4 . (41)
Simplifying, we find $\lambda\zeta^{2}>\frac{1}{4}$, which is satisfied by the initial assumption. ∎
See 3.7
Proof.
We begin the proof with the following observation: $\lambda$-JME and CS introduce additive noise to the first and second moments. Therefore, for the Frobenius approximation error, it is sufficient to compare just the variances of the noise introduced by those methods. We use an instance of a Gaussian mechanism, and the privacy guarantees are more appropriately characterized by the notion of Gaussian privacy. For the sake of the proof, we assume that $(\epsilon,\delta)$-DP is equivalent to $\mu$-GDP for a specific choice of $\mu$. The sensitivity of $\lambda$-JME depends on the dimensionality. Here, we assume $d>1$ to address the hardest case, as $d=1$ follows trivially.
We aim to show that Joint Moment Estimation (JME) introduces less noise to the $(x\otimes x)$ component than CS under the same privacy budget $\mu$-GDP. Specifically, we compare the variances in the second coordinate under the assumption that the noise variances in the first coordinate are equal.
The combined vector of $x_{i}$ and $x_{i}\otimes x_{i}$ can be represented as:
$(x_{i},\sqrt{\tau}x_{i}\otimes x_{i}),$ (42)
where $\tau>0$ is a scaling parameter. If the input vectors $x_{i}$ are bounded by $\|x_{i}\|_{2}\leq\zeta$, the resulting vector will have $l_{2}$ norm bounded by $\sqrt{\zeta^{2}+\tau\zeta^{4}}$. The squared sensitivity of the vector is therefore: $4\zeta^{2}(1+\tau\zeta^{2})$. This results in the following variances for the first and second coordinates under noise addition:
$\left(\frac{4\zeta^{2}(1+\tau\zeta^{2})}{\mu^{2}},\frac{4\zeta^{2}(1+\tau\zeta^{2})}{\tau\mu^{2}}\right).$ (43)
For JME, we analyze the variances and obtain the pair of variances as:
$\left(\frac{\zeta^{2}(2+2\zeta^{2}\lambda+\frac{1}{2\zeta^{2}\lambda})}{\mu^{2}},\frac{\zeta^{2}(2+2\zeta^{2}\lambda+\frac{1}{2\zeta^{2}\lambda})}{\lambda\mu^{2}}\right),$ (44)
where we assume $\zeta^{2}\lambda\geq\frac{1}{2}$.
To compare the noise introduced to the second component, we aim to show:
$\frac{\zeta^{2}(2+2\zeta^{2}\lambda+\frac{1}{2\zeta^{2}\lambda})}{\lambda\mu^{2}}<\frac{4\zeta^{2}(1+\tau\zeta^{2})}{\tau\mu^{2}},$ (45)
under the condition that the variances in the first component are equal:
$\frac{\zeta^{2}(2+2\zeta^{2}\lambda+\frac{1}{2\zeta^{2}\lambda})}{\mu^{2}}=\frac{4\zeta^{2}(1+\tau\zeta^{2})}{\mu^{2}}.$ (46)
Given the equality, inequality is equivalent to $\lambda>\tau$. Simplifying the equality in the first component gives:
$2+2\zeta^{2}\lambda+\frac{1}{2\zeta^{2}\lambda}=4+4\tau\zeta^{2}.$ (47)
Rearranging terms, we isolate $\frac{1}{2\zeta^{2}\lambda}$:
$\frac{1}{2\zeta^{2}\lambda}=2+2(2\tau-\lambda)\zeta^{2}.$ (48)
From the assumption $\zeta^{2}\lambda\geq\frac{1}{2}$, we know that $\frac{1}{2\zeta^{2}\lambda}<1$. For this inequality to hold, the term $2+2(2\tau-\lambda)\zeta^{2}$ must also be less than 1. Therefore, $02\tau-\lambda<$ 0, therefore $\lambda>\tau$.
Thus, we conclude that for the same variance in the first component, JME introduces less noise variance to the second component compared to CS. ∎
See 3.8
Proof.
We aim to evaluate the expected squared Frobenius norm of the error:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(X+C_{1}^{-1}Z_{1}))-A_{2}(X\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}$
$\quad=2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{\mathcal{S}_{1}}+2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\!\bullet\!}X)\rangle_{\text{F}}}_{\mathcal{S}_{2}}+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{\mathcal{S}_{3}}.$
We compute those terms separately.
Step 1. Bound $\mathcal{S}_{1}$.
$\begin{split}\mathcal{S}_{1}&=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}\\ &=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,i}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}\\ &=\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{1}}X_{t_{1},i}X_{t_{2},i}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\mathbb{E}(Z_{1})_{r,j}^{2}\\ &=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}X_{t_{1},i}X_{t_{2},i}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\\ &=d\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\sum\limits_{t_{1},t_{2}}^{n}\langle(A_{2}^{\top})_{t_{1}},(A_{2}^{\top})_{t_{2}}\rangle\langle X_{t_{1}},X_{t_{2}}\rangle\langle(C_{1}^{-1})_{t_{1}},(C_{1}^{-1})_{t_{2}}\rangle\\ &=d\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T}),\end{split}$ (49)
where $Q=(C_{1}C_{1}^{T})^{-1}$.
Step 2. Bound $\mathcal{S}_{2}$.
$\begin{split}\mathcal{S}_{2}&=\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\!\bullet\!}X)\rangle_{\text{F}}\\ &=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,i}(C_{1}^{-1}Z_{1})_{t,j}\right)\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}(C_{1}^{-1}Z_{1})_{t,i}\right)\\ &=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}X_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(A_{2})_{k,t_{2}}X_{t_{2},j}(C_{1}^{-1}Z_{1})_{t_{2},i}\\ &=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}X_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{1},j}(A_{2})_{k,t_{2}}X_{t_{2},j}(C_{1}^{-1}Z_{1})_{t_{2},j}\\ &=\frac{1}{d}\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\frac{S_{1}}{d}\end{split}$ (50)
Step 3. Bounding $\mathcal{S}_{3}$ We expand this term:
$\begin{split}\mathcal{S}_{3}&=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}\\ &=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})_{t,i}(C_{1}^{-1}Z_{1})_{t,j}\right)^{2}\\ &=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}\\ &=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}\sum\limits_{i,j}^{d}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}\\ &=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}\sum\limits_{i,j}^{d}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}\\ &=\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}\sum\limits_{i,j}^{d}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}.\end{split}$ (51)
We now deal with two cases. When $i=j$, we compute:
$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}&=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}\\ &\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (52)
On the other hand, when $i\neq j$, then
$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}&=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\right)^{2}=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (53)
Using equation (52) and equation (53) in equation (51), we get
$\mathcal{S}_{3}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})+d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q)),$
where $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}^{\top}(Q)$.
Adding equation (49) to equation (51), we obtain:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+2(d+1)\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
$\quad+d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
Bias Correction.
The expectation of $A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))]_{k,i,j}$ introduces a bias:
$[\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))]_{k,i,j}$ $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,i}\right)\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)$ (54)
$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
The Frobenius norm of this bias is:
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
If we subtract this bias from the estimate, it will increase the error by the aforementioned quantity due to the Frobenius norm of the bias but will decrease the error by two scalar products with the $A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))$ term:
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction. ∎
We collect some useful proposition that would be useful for our analysis.
Proposition C.2.
Let $V$ and $X$ be a given fixed matrix and $Z$ be a Gaussian matrix of appropriate dimension. Then
$\mathbb{E}_{Z}\|VZ\operatorname{\!\bullet\!}VZ\|_{\text{F}}^{2}=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.$
Proof.
Recalling the Face-Splitting product, we have
$\mathbb{E}_{Z}\|VZ\operatorname{\!\bullet\!}VZ\|_{\text{F}}^{2}$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ)^{2}_{k,i}(VZ)_{k,j}^{2}=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}Z_{t,i}\right)^{2}\left(\sum\limits_{t=1}^{n}V_{k,t}Z_{t,j}\right)^{2}$ (55)
$=3\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{i=1}^{d}\sum\limits_{t=1}^{n}V_{k,t}^{4}+\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{i\neq j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}+3\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{i=1}^{d}\sum\limits_{j\neq\ell}^{n}V_{k,j}^{2}V_{k,\ell}^{2}$
$=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2},$
This completes the proof of the proposition. ∎
Proposition C.3.
Let $V$ and $X$ be a given fixed matrix and $Z$ be a Gaussian matrix of appropriate dimension. Then
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\!\bullet\!}VX\|_{\text{F}}^{2}.$
Proof.
The result follows using the following calculation:
$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ_{1})_{k,i}(VX)_{k,j}(VZ_{1})_{k,j}(VX)_{k,i}=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\left(\sum\limits_{i=1}^{d}(VZ_{1})_{k,i}(VX)_{k,i}\right)^{2}$
$=\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\!\bullet\!}VX\|_{\text{F}}^{2}$
completing the proof. ∎
See 4.1
Proof.
We first recall that, if $Z\sim\mathcal{N}(\mu,\Sigma)$, then for any matrix $A$, we have $AZ\sim\mathcal{N}(A\mu,A\Sigma A^{\top})$. This implies that, when $\Sigma=\mathbb{I}$ and $\mu=0$, we have
$\mathbb{E}_{Z}\|AZ\|_{\text{F}}^{2}=\|A\|_{\text{F}}^{2}.$
Recall that $\widehat{\mu}=V(X+Z_{1})$ is a running mean and $\widehat{\Sigma}=V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))$ is a running covariance matrix, with independent noise $Z_{1},Z_{2}\in\mathcal{N}(0,\sigma^{2})^{n\times d^{2}}$, the clipping norm $\zeta=1$, and $\lambda=\lambda^{*}=c_{d}^{-1}$ as defined in equation (22). Using the associativity of Face-splitting product and the Pythogorean theorem, we have
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}$
$\quad=\mathbb{E}_{Z_{1},Z_{2}}\|\lambda^{-1/2}VZ_{2}-(VZ_{1})\operatorname{\!\bullet\!}(VX)-(VX)\operatorname{\!\bullet\!}(VZ_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))\|_{\text{F}}^{2}$
$\quad=\mathbb{E}_{Z_{2}}\|\lambda^{-1/2}VZ_{2}\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle$
$\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}}+2{\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}$
Using Proposition C.2 and Proposition C.3, we therefore have
$\begin{split}&\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2\left(1+{1\over d}\right){\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.\end{split}$ (56)
Now using the properties of $V$, we have
$\begin{split}\|V\|_{\text{F}}^{2}=\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}=\sum\limits_{k=1}^{n}\frac{1}{k}=H_{n,1}\quad\text{and}\quad\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=H_{n,2}.\end{split}$ (57)
Using equation (57) in equation (LABEL:eq:simplying), we get
$\begin{split}&\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=2\left(1+{1\over d}\right){\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2\sigma^{2}d^{2}\sum\limits_{k=1}^{n}\frac{1}{k}\\ &\quad=2\left(1+{1\over d}\right){\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+d(d+2)\sigma^{4}H_{n,2}+2\sigma^{2}d^{2}H_{n,1}.\end{split}$ (58)
Let $H_{n,c}$ denote the generalized Harmonic sum, i.e., $H_{n,c}=\sum_{i=1}^{n}i^{-c}$. Therefore, we have
$\begin{split}&\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=2\left(1+{1\over d}\right)\underbrace{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}_{S(X)}+\sigma^{4}d(d+2)H_{n,2}+c_{d}\sigma^{2}d^{2}H_{n,1}.\end{split}$ (59)
Therefore to estimate $\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$, it suffices to estimate $\sup_{X\in\mathcal{X}}S(X)$. We do it as follows:
$\sup_{X\in\mathcal{X}}S(X)$ $=\sup_{X\in\mathcal{X}}\mathbb{E}_{Z_{1}}\|VZ_{1}\operatorname{\!\bullet\!}VX\|_{\text{F}}^{2}=d\sigma^{2}\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{3}}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{k}X_{t,j}\right)^{2}$ (60)
$=d\sigma^{2}\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{3}}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}\rangle=d\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}=d\sigma^{2}H_{n,1},$
Plugging equation (60) in equation (LABEL:eq:simplifiedend), we get the bound for the biased estimate:
$\sigma^{2}(c_{d}d^{2}+2d+2)H_{n,1}+\sigma^{4}d(d+2)H_{n,2}.$ (61)
Then we need to determine the bias term, all the first-order terms will result in 00 as the expectation is over a zero mean distribution, so the only term that introduces the bias is
$(\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))_{k,i,j}=\mathbb{E}_{Z_{1}}(VZ_{1})_{k,i}(VZ_{1})_{k,j}=\sigma^{2}\delta_{i=j}\sum\limits_{t=1}^{n}V_{k,t}^{2},$ (62)
where $\delta_{i=j}=\begin{cases}1&i=j\\ 0&\text{otherwise}\end{cases}$ is the Dirac-delta function.
Then the unbiased approximation error is
$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)+\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$
We have already computed the first three terms inside the expectation. We now compute the bias term error:
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
Bias reduction procedure decreases the expected error of (61) by
$-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
Resulting in the final error of:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=\sigma^{2}(c_{d}d^{2}+2d+2)\sum\limits_{k=1}^{n}\frac{1}{k}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
$=\sigma^{2}(c_{d}d^{2}+2d+2)H_{n,1}+d(d+1)\sigma^{4}H_{n,2}.$
This completes the proof of the theorem. ∎
See 4.2
Proof.
Let us denote the covariance matrix estimated via Post-Processing (PP) without bias correction as $\widehat{\Sigma}^{b}_{\text{PP}}$. Then, the approximation error has the following form:
$\begin{split}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}=&\mathbb{E}_{Z_{1}}\|V((X+Z_{1})\operatorname{\!\bullet\!}(X+Z_{1}))-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})+V(Z_{1}\operatorname{\!\bullet\!}X)+V(X\operatorname{\!\bullet\!}Z_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VX)-(VX)\operatorname{\!\bullet\!}(VZ_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))\|_{\text{F}}^{2}\\ &\quad=\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}}_{A_{1}}-2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}Z_{1}),(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{2}}+2\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}}_{A_{3}}\\ &\hskip 56.9055pt+2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),V(X\operatorname{\!\bullet\!}Z_{1})\rangle}_{A_{4}}-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{5}}\\ &\hskip 56.9055pt-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VZ_{1})\operatorname{\!\bullet\!}(VX)\rangle}_{A_{6}}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle\end{split}$ (63)
The last three terms in the above expression evaluates to equation (LABEL:eq:simplifiedend). Therefore, in what follows, we bound $A_{1}$ to $A_{6}$.
Bounding $A_{1}$:
$\begin{split}A_{1}&=\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}(Z_{1})_{t,j}\right)^{2}\\ &=3\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t=1}^{n}V_{k,t}^{2}+\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1}\neq t_{2}}^{n}V_{k,t_{1}}V_{k,t_{2}}+\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{i\neq j}^{d}\sum\limits_{t=1}^{n}V_{k,t}^{2}\\ &=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}\\ \end{split}$ (64)
Bounding $A_{2}$:
$\begin{split}A_{2}&=\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}Z_{1}),(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle\\ &=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}(Z_{1})_{t,j}\right)\left(\sum\limits_{t}V_{k,t}(Z_{1})_{t,i}\right)\left(\sum\limits_{t}V_{k,t}(Z_{1})_{t,j}\right)\\ &=3\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t=1}^{n}V_{k,t}^{3}+\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{i\neq j}^{d}\sum\limits_{t=1}^{n}V_{k,t}^{3}+\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1}\neq t_{2}}^{n}V_{k,t_{1}}V_{k,t_{2}}^{2}\\ &=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)\left(\sum\limits_{t=1}^{n}V^{2}_{k,t}\right)\end{split}$ (65)
Bounding $A_{3}$:
$A_{3}$ $=\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}X_{t,j}\right)^{2}=d\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}$ (66)
Bounding $A_{4}$:
$\begin{split}A_{4}&=\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),V(X\operatorname{\!\bullet\!}Z_{1})\rangle\\ &=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,j}X_{t,i}\right)=\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}\end{split}$ (67)
Bounding $A_{5}$:
$\begin{split}A_{5}&=\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle\\ &=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,i}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,j}\right)\\ &=\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)\end{split}$ (68)
Bounding $A_{6}$:
$\begin{split}A_{6}&=\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VZ_{1})\operatorname{\!\bullet\!}(VX)\rangle\\ &=\mathbb{E}_{Z_{1}}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}(Z_{1})_{t,i}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)\\ &=d\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)\end{split}$ (69)
Plugging equation (64) to equation (69) in equation (63), we get
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
$\quad-2d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)\left(\sum\limits_{t=1}^{n}V^{2}_{k,t}\right)$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$
Recalling that the matrix $V$ is the averaging workload matrix $V=(a^{t}_{i})$ with $a^{t}_{i}=\frac{1}{t}$ for $1\leq i\leq t$ and $a^{t}_{i}=0$ otherwise, we get
$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}-2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{k}\frac{1}{k^{2}}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{k}\frac{1}{k^{2}}X_{t,j}\right)\left(\sum\limits_{t=1}^{k}\frac{1}{k}X_{t,j}\right)$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{k}\frac{1}{k}X_{t,j}\right)^{2}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
$=d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\underbrace{\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{1},:}\rangle\right]}_{T_{n}(X)}$
Since every term except $T_{n}(X)$ is independent of $X$, to compute both upper and lower bounds on $\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$, it suffices to bound the supremum of the inner difference over $X$:
$T_{n}:=\sup_{X\in\mathcal{X}}T_{n}(X)\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\underbrace{\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}}_{=\left\|\sum_{t}X_{t}\right\|^{2}_{2}\geq 0}\rangle\right].$
Since the double sum is non-negative, this leads to a trivial upper bound $T_{n}\leq\sum\limits_{k=1}^{n}\frac{1}{k}=H_{n,1}$. For the lower bound, consider $X_{i}=(-1)^{i}e_{1}$. In this case, $\|X_{i}\|_{2}=1$, but $\left\|\sum_{t}X_{t}\right\|^{2}_{2}\leq 1$, so $\sum\limits_{k=1}^{n}\left(\frac{1}{k}-\frac{1}{k^{3}}\right)\leq T_{n}$. Therefore,
$H_{n,1}-H_{n,3}\leq T_{n}\leq H_{n,1}.$
Therefore, we have the following bounds for the approximation error without a bias correction:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}$
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}-2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k^{3}},$
Now we will determine the bias term; let $\delta_{i=j}$ denote the Dirac-delta function, then the bias of $\widehat{\Sigma}_{\text{PP}}$ is
$(\mathbb{E}_{Z_{1}}(V(Z_{1}\operatorname{\!\bullet\!}Z_{1})-\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))_{k,i,j}=\sigma^{2}\delta_{i=j}\sum\limits_{t=1}^{n}V_{k,t}-V_{k,t}^{2}$ (70)
We also have the following set of equalities when $Z_{1}$ is a Gaussian matrix.
$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}=dn\sigma^{4}$
$\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle_{\text{F}}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
To remove the bias from the error we would need to add the following terms:
$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}-2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle$
$\quad-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle_{\text{F}}+4\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle$
$\quad-2\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}Z_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle_{\text{F}}$
$=-\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}-\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}+2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle$
$=-d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}+2d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)$
$=-dn\sigma^{4}-d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}=-dn\sigma^{4}-d\sigma^{4}H_{n,2}+2d\sigma^{4}H_{n,1}.$
Combining everything together we obtain:
$\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$
Therefore,
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}\quad\text{and}$
$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
which concludes the proof. ∎
See A.5
Proof.
We begin by selecting indices $i$ and $j$ such that the corresponding components $x_{i},x_{j}$ from $x$ and $y_{i},y_{j}$ from $y$ satisfy $(x_{i}^{2}-y_{i}^{2})(x_{j}^{2}-y_{j}^{2})\geq 0$. We can always find such indices because, by the pigeonhole principle for $d\geq 3$, there will be pairs where either both $x_{i}^{2}\geq y_{i}^{2}$ and $x_{j}^{2}\geq y_{j}^{2}$, or both $x_{i}^{2}\leq y_{i}^{2}$ and $x_{j}^{2}\leq y_{j}^{2}$. We can compare the impact of these components on the sum with the values $\sqrt{x_{i}^{2}+x_{j}^{2}}$ and $-\sqrt{y_{i}^{2}+y_{j}^{2}}$, which correspond to vectors in a lower dimension. Consider the difference in the objective function and $f(x_{,}i,y_{i})$ defined below:
$f(x_{i},y_{i})$ $:=\left(\sqrt{x_{i}^{2}+x_{j}^{2}}+\sqrt{y_{i}^{2}+y_{j}^{2}}\right)^{2}+\lambda\left(x_{i}^{2}+x_{j}^{2}-y_{i}^{2}-y_{j}^{2}\right)^{2}$
$g(x_{i},y_{i})$ $:=(x_{i}-y_{i})^{2}+(x_{j}-y_{j})^{2}+\lambda(x_{i}^{2}-y_{i}^{2})^{2}+\lambda(x_{j}^{2}-y_{j}^{2})^{2}.$
Note that $f(x_{i},y_{i})$ is the objective function. Now, after algebraic manipulation, we get
$f(x_{i},y_{i})-g(x_{i},y_{i})$ $=2\sqrt{x_{i}^{2}+x_{j}^{2}}\sqrt{y_{i}^{2}+y_{j}^{2}}+\lambda(x_{i}^{2}-y_{i}^{2})(x_{j}^{2}-y_{j}^{2})+2x_{i}y_{i}+2x_{j}y_{j}$ (71)
$\geq\lambda(x_{i}^{2}-y_{i}^{2})(x_{j}^{2}-y_{j}^{2})\geq 0,$ (72)
where the first inequality follows from the Cauchy-Schwarz inequality and the second inequality is from the assumption that $(x_{i}^{2}-y_{i}^{2})(x_{j}^{2}-y_{j}^{2})\geq 0$.
For this lemma, $d=2$ is indeed a special case since it is possible to find $x_{1}^{2}>y_{1}^{2}$ and $x_{2}^{2}<y_{2}^{2}$, for which the dimension reduction argument would not work. ∎
See A.4
Proof.
First, we consider the effect of the signs of the components of $x$ and $y$. Multiplying both $x_{i}$ and $y_{i}$ by $-1$ does not change the value of the expression. If $x_{i}$ and $y_{i}$ have the same sign, then by flipping the sign of one of them, we increase the difference $\|x-y\|_{2}$, while $\|x\circ x-y\circ y\|_{2}$ remains unchanged. Therefore, without loss of generality, we can assume $x=(x_{1},x_{2})$ and $y=(-y_{1},-y_{2})$ for positive $x_{i}$ and $y_{i}$. Next, note that for a fixed difference $\|x-y\|_{2}$, the functional increases as the sum $x_{1}+y_{1}$ or $x_{2}+y_{2}$ increases. Thus, we can assume $\|x\|_{2}=1$, so $x_{2}=\sqrt{1-x_{1}^{2}}$. The norm of $y$, however, can be different. Therefore, we look for a solution of the form $x=(x_{1},\sqrt{1-x_{1}^{2}})$ and $y=(-y_{1},-y_{2})$.
We now consider the functional in the statement of the lemma in the following form:
$\mathcal{L}_{\lambda}(x_{1},y_{1},y_{2})=(x_{1}+y_{1})^{2}+\left(\sqrt{1-x_{1}^{2}}+y_{2}\right)^{2}+\lambda\left(x_{1}^{2}-y_{1}^{2}\right)^{2}+\lambda\left(1-x_{1}^{2}-y_{2}^{2}\right)^{2}.$ (73)
We now aim to prove that $\sup\mathcal{L}_{\lambda}(x_{1},y_{1},y_{2})=r_{2}^{\text{diag}}(\lambda)$ in the constrained domain $y_{1}\geq 0$, $y_{2}\geq 0$, $y_{1}^{2}+y_{2}^{2}\leq 1$, $0\leq x_{1}\leq 1$. This analysis involves considering up to twenty-four different scenarios for optimization with boundaries, which, due to symmetry, can be reduced to five distinct cases.
Case I: $y_{1}^{2}+y_{2}^{2}=1$, $y_{1}>0$, $y_{2}>0$.
We can compute $y_{2}=\sqrt{1-y_{1}^{2}}$, then the optimization functional is
$\mathcal{L}_{\lambda}\left(x_{1},y_{1},\sqrt{1-y_{1}^{2}}\right)$ $=(x_{1}+y_{1})^{2}+\left(\sqrt{1-x_{1}^{2}}+\sqrt{1-y_{1}^{2}}\right)^{2}+\lambda\left(x_{1}^{2}-y_{1}^{2}\right)^{2}+\lambda\left(y_{1}^{2}-x_{1}^{2}\right)^{2}$ (74)
$=2+2x_{1}y_{1}+2\sqrt{1-x_{1}^{2}}\sqrt{1-y_{1}^{2}}+2\lambda\left(x_{1}^{2}-y_{1}^{2}\right)^{2}.$ (75)
The cases where $x_{1}=0$ or $x_{1}=1$ with $\|y\|=1$ are equivalent to the scenario where $\|x\|=1$ and $y_{1}=0$, or $y_{1}=1$, which we will consider later. For now, we proceed by computing the derivatives with respect to $y_{1}$ and $x_{1}$:
$\begin{split}\frac{\partial\mathcal{L}_{\lambda}\left(x_{1},y_{1},\sqrt{1-y_{1}^{2}}\right)}{\partial y_{1}}=2x_{1}-\frac{2y_{1}\sqrt{1-x_{1}^{2}}}{\sqrt{1-y_{1}^{2}}}+4y_{1}\lambda\left(y_{1}^{2}-x_{1}^{2}\right)=0,\\ \frac{\partial\mathcal{L}_{\lambda}\left(x_{1},y_{1},\sqrt{1-y_{1}^{2}}\right)}{\partial x_{1}}=2y_{1}-\frac{2x_{1}\sqrt{1-y_{1}^{2}}}{\sqrt{1-x_{1}^{2}}}+4x_{1}\lambda\left(x_{1}^{2}-y_{1}^{2}\right)=0.\end{split}$ (76)
We transform this system by summing and subtracting the equalities:
$\begin{split}\left(x_{1}+y_{1}\right)-\frac{y_{1}-y_{1}x_{1}^{2}+x_{1}-x_{1}y_{1}^{2}}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}+4\lambda\left(x_{1}^{3}+y_{1}^{3}\right)-4\lambda y_{1}x_{1}\left(y_{1}+x_{1}\right)=0,\\ \left(x_{1}-y_{1}\right)-\frac{y_{1}-y_{1}x_{1}^{2}-x_{1}+x_{1}y_{1}^{2}}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}+4\lambda\left(y_{1}^{3}-x_{1}^{3}\right)-4\lambda y_{1}x_{1}\left(x_{1}-y_{1}\right)=0.\end{split}$ (77)
$x_{1}+y_{1}=0$ is not a solution under the constraints we are solving. However, $x_{1}=y_{1}$ is a solution that gives $\mathcal{L}_{\lambda}(x_{1},x_{1},\sqrt{1-x_{1}^{2}})=4$ for any $\lambda$. From now on, consider $y_{1}\neq x_{1}$; then we can divide by the difference, leading to the system:
$\begin{split}1-\frac{1-x_{1}y_{1}}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}+4\lambda\left(x_{1}-y_{1}\right)^{2}=0,\\ 1+\frac{1+x_{1}y_{1}}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}-4\lambda\left(x_{1}+y_{1}\right)^{2}=0.\end{split}$ (78)
We consider again the sum and difference to get:
$\begin{split}2+\frac{2x_{1}y_{1}}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}-16\lambda x_{1}y_{1}=0,\\ -\frac{2}{\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}+8\lambda\left(x_{1}^{2}+y_{1}^{2}\right)=0.\end{split}$ (79)
We solve it for $\lambda$ to get the following equation:
$\frac{1}{4\left(x_{1}^{2}+y_{1}^{2}\right)\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}$ $=\frac{1}{8x_{1}y_{1}}+\frac{1}{8\sqrt{1-y_{1}^{2}}\sqrt{1-x_{1}^{2}}}.$ (80)
We transform it into the form
$\frac{2x_{1}y_{1}}{x_{1}^{2}+y_{1}^{2}}-x_{1}y_{1}=\sqrt{1-y_{1}^{2}-x_{1}^{2}+x_{1}^{2}y_{1}^{2}}.$ (81)
By squaring both sides and subtracting $x_{1}^{2}y_{1}^{2}$, we obtain
$\frac{4x_{1}^{2}y_{1}^{2}}{(x_{1}^{2}+y_{1}^{2})^{2}}(1-x_{1}^{2}-y_{1}^{2})=1-x_{1}^{2}-y_{1}^{2}.$ (82)
So, either $x_{1}^{2}+y_{1}^{2}=1$, which implies $y_{1}=\sqrt{1-x_{1}^{2}}$, or $x_{1}^{2}+y_{1}^{2}=2x_{1}y_{1}$, which implies $x_{1}=y_{1}$, which we have already discussed. We have found another potentially optimal point $y_{1}=\sqrt{1-x_{1}^{2}}$, which we will further investigate. We now consider it as a function of one variable $x_{1}$:
$\mathcal{L}_{\lambda}\left(x_{1},\sqrt{1-x_{1}^{2}},x_{1}\right)=2\left(x_{1}+\sqrt{1-x_{1}^{2}}\right)^{2}+2\lambda\left(1-2x_{1}^{2}\right)^{2}.$ (83)
Then the optimum is either $x_{1}=0$, $x_{1}=1$ with the value $4$, or when the derivative is 00:
$\frac{\partial\mathcal{L}_{\lambda}\left(x_{1},\sqrt{1-x_{1}^{2}},x_{1}\right)}{\partial x_{1}}$ $=4\sqrt{1-x_{1}^{2}}-\frac{4x_{1}^{2}}{\sqrt{1-x_{1}^{2}}}-16\lambda x_{1}(1-2x_{1}^{2})$ (84)
$=\frac{4(1-2x_{1}^{2})\left(1-4\lambda x_{1}\sqrt{1-x_{1}^{2}}\right)}{\sqrt{1-x_{1}^{2}}}=0.$ (85)
The first term gives $x_{1}=\frac{1}{\sqrt{2}}$, which results in $\mathcal{L}_{\lambda}=4$; otherwise, $14\lambda x_{1}\sqrt{1-x_{1}^{2}}=$ 1, which we can solve by first denoting $x_{1}^{2}$ as a new variable in the equation $-x_{1}^{4}+x_{1}^{2}-\frac{1}{16\lambda^{2}}=0$. Using the quadratic formula, we find the optimal value $x_{1}^{*}$, which is given by:
$x_{1}^{*}=\sqrt{\frac{1}{2}+\frac{1}{2}\sqrt{1-\frac{1}{4\lambda^{2}}}},\quad\sqrt{1-(x_{1}^{*})^{2}}=\sqrt{\frac{1}{2}-\frac{1}{2}\sqrt{1-\frac{1}{4\lambda^{2}}}}.$ (86)
This solution exists only when $\lambda\geq\frac{1}{2}$. Substituting this root back into the function gives us
$\mathcal{L}_{\lambda}\left(x_{1}^{*},\sqrt{1-(x_{1}^{*})^{2}},x_{1}^{*}\right)=2+\underbrace{4x_{1}^{*}\sqrt{1-(x_{1}^{*})^{2}}}_{=\frac{1}{\lambda}}+2\lambda+\underbrace{8\lambda((x_{1}^{*})^{4}-(x_{1}^{*})^{2})}_{=-\frac{1}{2\lambda}}=2+2\lambda+\frac{1}{2\lambda}\geq 4,$ (87)
which constitutes the function $r_{2}^{\text{diag}}(\lambda)$.
Case II: $y_{1}^{2}+y_{2}^{2}<1$, $y_{1},y_{2}>0$ (Interior).
For the optimum, it is important that $\frac{\partial\mathcal{L}_{\lambda}}{\partial y_{1}}=0$ and $\frac{\partial\mathcal{L}_{\lambda}}{\partial y_{2}}=0$. But the necessary condition for the maximum would be that the Hessian is negative semi-definite: $\frac{\partial^{2}\mathcal{L}_{\lambda}}{\partial y_{1}^{2}}\leq 0$ and $\frac{\partial^{2}\mathcal{L}_{\lambda}}{\partial y_{2}^{2}}\leq 0$. We can compute the second derivatives explicitly:
$\frac{\partial^{2}\mathcal{L}_{\lambda}}{\partial y_{1}^{2}}=2+12\lambda y_{1}^{2}-4\lambda x_{1}^{2}\leq 0\Rightarrow y_{1}^{2}\leq\frac{x_{1}^{2}}{3}.$ (88)
Analogously, we can derive that $y_{2}^{2}\leq\frac{1-x_{1}^{2}}{3}$. Then substituting this into the functional, we get:
$\mathcal{L}_{\lambda}(x_{1},y_{1},y_{2})$ $=(x_{1}+y_{1})^{2}+\left(\sqrt{1-x_{1}^{2}}+y_{2}\right)^{2}+\lambda(x_{1}^{2}-y_{1}^{2})^{2}+\lambda\left(1-x_{1}^{2}-y_{2}^{2}\right)^{2}$ (89)
$\leq\left(1+\frac{1}{\sqrt{3}}\right)^{2}+\lambda+\frac{\lambda}{9}<r_{2}(\lambda).$ (90)
Case III: $y_{1}=0$, $10<y_{2}<$ 1.
$\mathcal{L}_{\lambda}(x_{1},y_{1},y_{2})$ $=x_{1}^{2}+\left(\sqrt{1-x_{1}^{2}}+y_{2}\right)^{2}+\lambda x_{1}^{4}+\lambda\left(1-x_{1}^{2}-y_{2}^{2}\right)^{2}.$ (91)
As a continuous function of $y_{2}$ on an open domain, it can reach a maximum only when the second derivative is non-positive, which leads to the same condition as in the previous case: $y_{2}^{2}\leq\frac{1-x_{1}^{2}}{3}$. Since $y_{1}=0\leq\frac{x_{1}^{2}}{3}$, this leads to the same conclusion.
Case IV: $y_{1}=0$, $y_{2}=0$.
$\begin{split}\mathcal{L}_{\lambda}(x_{1},0,0)&=x_{1}^{2}+1+\lambda x_{1}^{4}+\lambda(1-x_{1}^{2})^{2}\leq x_{1}^{2}+\lambda x_{1}^{4}+\lambda(1-x_{1}^{2})^{2}+1\leq 1+\lambda<r_{2}^{\text{diag}}(\lambda).\end{split}$ (92)
Case V: $y_{1}=1$, $y_{2}=0$.
$\mathcal{L}_{\lambda}(x_{1},1,0)$ $=(x_{1}+1)^{2}+1-x_{1}^{2}+\lambda(1-x_{1}^{2})^{2}+\lambda(1-x_{1}^{2})^{2}=2+2x_{1}+2\lambda(1-x_{1}^{2})^{2}.$ (93)
First, for $\lambda\leq\frac{1}{2}$, we have:
$\mathcal{L}_{\lambda}(x_{1},1,0)$ $\leq 2+2x_{1}+1-x_{1}^{2}\leq 4.$ (94)
For $\lambda>\frac{1}{2}$, we have the following inequality:
$\mathcal{L}_{\lambda}(x_{1},1,0)$ $\leq 2+2x_{1}+2\lambda(1-x_{1}^{2})=2+2\lambda+2x_{1}-2\lambda x_{1}^{2}\leq 2+2\lambda+\frac{1}{2\lambda}.$ (95)
This concludes the proof. ∎
Lemma C.4 (Expected Second Moment Error with PP).
Given the private estimation of the first moment $A_{1}(X+C_{1}^{-1}Z_{1})$, where $A_{1}=B_{1}C_{1}$, and the second moment $A_{2}(X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1})$, where $A_{2}=B_{2}C_{2}$, with independent noise $Z_{1}\in\mathcal{N}(0,\|C_{1}\|_{1\to 2}^{2}\sigma^{2})^{n\times d}$, the clipping norm $\zeta=1$, $d>1$, the expected squared Frobenius norm of the estimation error for the second moment satisfies:
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}:$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$ (96)
$=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
$\quad+\underbrace{d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}_{\text{bias}}$
where $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}^{\top}(Q)$.
Proof.
We aim to evaluate the expected squared Frobenius norm of the error:
$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$
$\quad=4\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{S_{1}}+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{S_{2}}$
We compute those terms separately.
$S_{1}$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$
$=\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{1}}X_{t_{1},j}X_{t_{2},j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\mathbb{E}(Z_{1})_{r,j}^{2}$
$=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}X_{t_{1},j}X_{t_{2},j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}$
$=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\sum\limits_{t_{1},t_{2}}^{n}\langle(A_{2}^{\top})_{t_{1}},(A_{2}^{\top})_{t_{2}}\rangle\langle X_{t_{1}},X_{t_{2}}\rangle\langle(C_{1}^{-1})_{t_{1}},(C_{1}^{-1})_{t_{2}}\rangle$
$=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T}),$
where $Q=C_{1}^{-1}C_{1}^{-T}$.
$S_{2}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})^{2}_{t,j}\right)^{2}$
$=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}.$
First, we compute:
$\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}$ $=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
$\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
$=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}$
$\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}$
$=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.$
Plugging it back, we get
$\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})+2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q)),$
where $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}^{\top}(Q)$.
Adding these terms together we obtain:
$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{\top})$
$\quad+d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
Bias Correction.
The expectation of $A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))]_{k,j}$ introduces a bias:
$[\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))]_{k,j}$ $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$
$=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}$
$=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
The Frobenius norm of this bias is:
$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
If we subtract this bias from the estimate, it will increase the error by the aforementioned quantity due to the Frobenius norm of the bias but will decrease the error by two scalar products with the $A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))$ term:
$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction. ∎
Appendix D Algorithms
Algorithm 2 Differentially Private JME Adam
  Input: Initial model $\theta_{0}\in\mathbb{R}^{d}$, dataset $D$, batchsize $b$, matrices $C_{\beta_{1}},C_{\beta_{2}}\in\mathbb{R}^{n\times n}$, model loss $\ell(\theta,d)$, clipnorm $\zeta$, noise multiplier $\sigma_{\epsilon,\delta}\geq 0$, learning rate $\alpha>0$, and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\varepsilon=10^{-8}$.
  $m_{0}\leftarrow 0$   // first moment initialization.
  $v_{0}\leftarrow 0$   // second moment initialization.
  $\lambda,s_{\lambda}=\cdot\textit{Joint-sens}(C_{\beta_{1}},C_{\beta_{2}})$   // joint sensitivity
  $Z_{1},Z_{2}\sim N(0,\sigma_{\epsilon,\delta}^{2}s_{\lambda}^{2}I_{d})$   // noise generating
  for $i=1,2,\dots,n$ do
     $S_{i}\leftarrow\{d_{1},\dots,d_{m}\}\subseteq D$ select a data batch
     $g_{j}\leftarrow\nabla_{\theta}\ell(\theta_{i-1},d_{j}))\quad\text{for $j=1,\dots,m$}$
     $x_{i}\leftarrow\sum_{j=1}^{m}\min(1,\zeta/||g_{j}||)g_{j}$
     $\widehat{x}_{i}\leftarrow x_{i}+\zeta[C_{\beta_{1}}^{-1}Z_{1}]_{[i,\cdot]}$
     $\widehat{x^{2}}_{i}\leftarrow x_{i}^{2}+\lambda^{-1/2}\zeta[C_{\beta_{2}}^{-1}Z_{2}]_{[i,\cdot]}$
     $m_{i}\leftarrow m_{i-1}\beta_{1}+(1-\beta_{1})\widehat{x}_{i}$
     $v_{i}\leftarrow v_{i-1}\beta_{2}+(1-\beta_{2})\widehat{x^{2}}_{i}$
     $\widehat{m}_{i}=m_{i}/(1-\beta_{1}^{i})$   // bias-correction
     $\widehat{v}_{i}=v_{i}/(1-\beta_{2}^{i})$   // bias-correction
     $\theta_{i}\leftarrow\theta_{i-1}-\alpha\widehat{m}_{i}/(\sqrt{\widehat{v}_{i}}+\epsilon)$
  end for
  $\Theta=(\theta_{1},\dots,\theta_{n})$
Algorithm 3 $\lambda$-JME
0:  input stream vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms)  (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
  $s\leftarrow\zeta\|C_{1}\|_{1\to 2}r_{d}\left(\frac{\lambda\zeta^{2}\|C_{2}\|_{1\to 2}^{2}}{\|C_{1}\|_{1\to 2}^{2}}\right)^{1/2}$ // joint sensitivity
  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
[…]
  end for
Algorithm 4 $\alpha$-IME (IME with budget split parameter $\alpha\in(0,1)$)
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
0:  privacy trade-off $\alpha\in(0,1)$
  $(\epsilon_{1},\delta_{1})\leftarrow(\alpha\epsilon,\alpha\delta)$ // 1st moment private budget
  $(\epsilon_{2},\delta_{2})\leftarrow((1-\alpha)\epsilon,(1-\alpha)\delta)$ // 2nd moment private budget
  $\sigma_{1}\leftarrow$ noise strength for $(\epsilon_{1},\delta_{1})$-dp Gaussian mechanism
  $s_{1}\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma^{2}_{1}s^{2}_{1})\big{]}^{n\times d}$ // 1st moment noise
  $\sigma_{2}\leftarrow$ noise strength for $(\epsilon_{2},\delta_{2})$-dp Gaussian mechanism
  $s_{2}\leftarrow\sqrt{2}\zeta^{2}\|C_{2}\|_{1\to 2}$ // sensitivity of 2nd moment
  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma^{2}s_{2}^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
[…]
  end for
Algorithm 5 $\tau$-CS (CS with 2nd moment rescaling parameter $\tau>0$)
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with
0:  workload matrices $A=(a^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  input dimension $d$, bound on input norm $\zeta>0$
0:  privacy parameters $(\epsilon,\delta)$
0:  (optional) noise shaping matrix $C$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
  $s\leftarrow 2\zeta\sqrt{1+\tau\zeta^{2}}$ // sensitivity based on norm of concatenated data
  $Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d(d+1)}$ // noise matrix for concatenated data
  for $t=1,2,\dots,n$ do
     $\widetilde{x_{t}}=\big{(}\,x_{t},\sqrt{\tau}\textsf{vec}(x_{t}\otimes x_{t})\,\big{)}$
     $\widehat{\widetilde{x_{t}}}\leftarrow\widetilde{x_{t}}+[C^{-1}Z]_{[t,\cdot]}$
[…]
  end for
Algorithm 6 PP
0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
0:  noise shaping matrix $C_{1}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
0:  privacy parameters $(\epsilon,\delta)$
  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
  $Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
  for $t=1,2,\dots,n$ do
     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z]_{[t,\cdot]}$
     $b_{t}\leftarrow I_{d\times d}\times\sigma_{\epsilon,\delta}^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{k=1}^{n}(A_{2})_{t,k}(C_{1}C_{1}^{\top})^{-1}_{k,k}$ // bias term (optional)
     $\widehat{x_{t}\otimes x_{t}}\leftarrow\widehat{x_{t}}\otimes\widehat{x_{t}}-b_{t}$
[…]
  end for

### v2

@@
 Abstract
-We propose Joint Moment Estimation (JME), a method for continually and privately estimating both the first and second moments of data with reduced noise compared to naive approaches.
-JME uses the matrix mechanism and joint sensitivity analysis to allow the second-moment estimation with no additional privacy cost, thereby improving accuracy while maintaining privacy.
-We demonstrate JME’s effectiveness in two applications: estimating the running mean and covariance matrix for Gaussian density estimation and model training with DP-Adam on CIFAR-10.
-Machine Learning, ICML
+We propose Joint Moment Estimation (JME), a method for continually and privately estimating both the first and second moments of a data stream with reduced noise compared to naive approaches.
+JME supports the matrix mechanism and exploits a joint sensitivity analysis to identify a privacy regime in which the second-moment estimation incurs no additional privacy cost, thereby improving accuracy while maintaining privacy.
+We demonstrate JME’s effectiveness in two applications: estimating the running mean and covariance matrix for Gaussian density estimation and model training with DP-Adam.
 1 Introduction
-Estimating the first and second moments of data is a fundamental step in many machine learning algorithms, ranging from foundational methods, such as linear regression, principal component analysis, or Gaussian model fitting, to state-of-the-art neural network components, such as BatchNorm (Ioffe & Szegedy, 2015), and optimizers, such as Adam (Diederik & Ba, 2015).
-Often, these estimates must be computed and updated continuously, called the continual release setting (Dwork et al., 2010).
-For example, this is the case when the data arrives sequentially and intermediate results are needed without delay, such as for sequential optimization algorithms, such as Adam, or for real-time systems in healthcare, finance, or online recommendation systems.
+Estimating the first and second moments of data is a fundamental step in many machine learning algorithms, ranging from foundational methods, such as linear regression, principal component analysis, or Gaussian model fitting, to state-of-the-art neural network components, such as BatchNorm (pmlr-v37-ioffe15, ), and optimizers, such as Adam (kingma2014adam, ).
+Often, these estimates must be computed and updated continuously, called the continual release setting (dwork2010differential, ).
+For example, this is the case when the data arrives sequentially and intermediate results are needed without delay, such as for sequential optimization algorithms like Adam, or for real-time systems in healthcare, finance, or recommendation systems.
 In such scenarios, ensuring the privacy of sensitive data, such as user information or medical records, is essential.
 Differential Privacy (DP) is a widely established formal notion of privacy that ensures that the inclusion or exclusion of any single individual’s data has a limited impact on the output of an algorithm.
-Technically, DP masks sensitive information by the addition of suitably scaled noise, thereby creating a trade-off between privacy (formalized as a privacy budget) and utility (measured by the expected accuracy of the estimates).
+Technically, DP masks sensitive information by adding suitably scaled noise, thereby creating a trade-off between privacy (formalized as a privacy budget) and utility (measured by the expected accuracy of the estimates).
 This trade-off becomes particularly challenging when more than one quantity is meant to be estimated from the same private data, as is the case when estimating multiple data moments.
@@
 In this work, we introduce a new method, Joint Moment Estimation (JME), that is able to privately estimate the first and second moments of vector-valued data without suffering from this shortcoming.
-Algorithmically, JME relies on the recent matrix factorization (MF) mechanism (Li et al., 2015) for continual release DP to individually compute the first and the second moments of the data, thereby making it flexible to accommodate a variety of settings.
+Algorithmically, JME relies on the recent matrix factorization (MF) mechanism (li2015matrix, ) for continual release DP to individually compute the first and the second moments of the data, thereby making it flexible to accommodate a variety of settings.
 For example, besides the standard uniform sum or weighted average across data items, exponentially weighted averages or sliding-window estimates are readily possible.
 The key innovation of JME lies in our theoretical analysis of its properties.
-By jointly analyzing the sensitivity of the otherwise independent estimation processes, in combination with considering a carefully calibrated trade-off between them, we show that one can privately estimate the second-moment matrix without having to increase the amount of noise required to keep the first moment private.
+By jointly analyzing the sensitivity of the otherwise independent estimation processes, in combination with considering a carefully calibrated trade-off between them, we show that one can estimate privately the second-moment matrix without having to increase the amount of noise required to keep the first moment private.
 In this sense, we obtain privacy of the second moment for free.
@@
 Here, as well, we observe that JME achieves better optimization accuracy than baseline methods in high-privacy and small-batch-size regimes.
-method symbol coefficients
-prefix sum $E_{1}$ $\mathbbm{1}{\{i\leq t\}}$
-exponential (weight $\beta$) $E_{\beta}$ $\beta^{t-i}\mathbbm{1}\{i\leq t\}$
-standard average $V$ $\frac{1}{t}\mathbbm{1}\{i\leq t\}$
-sliding window (size $k$) $W_{k}$ $\frac{1}{k}\mathbbm{1}\{t-k\,<\,i\,\leq\,t\}$
-Table 1: Common workload matrices for moment estimation in (1).
 2 Background
-We study the problem of continually estimating the pair of weighted sums of the first and second moments in a differentially private manner.
-Namely, consider a sequence of $d$-dimensional vectors $x_{1},\dots,x_{n}\in\mathcal{X}$, where $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ for some fixed constant $\zeta>0$.
-At each step $t$, we aim to estimate the following pair of sums in a differentially private way:
+We study the problem of differentially private continually estimation of the pair of weighted sums of the first and second moments, see Section 6 for a discussion of related works.
+Specifically, consider a sequence of $d$-dimensional vectors $x_{1},\dots,x_{n}\in\mathcal{X}$, where $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;x_{i}\in\mathbb{R}^{d}\ \wedge\ \max_{i}\|x_{i}\|_{2}\leq\zeta\}$ for some fixed constant $\zeta>0$.
+At each step $t$, we aim to privately estimate the following pair of sums:
 $Y_{t}$ $=\sum\limits_{i=1}^{t}a^{t}_{i}x_{i}\in\mathbb{R}^{d}\ \ \text{and}\ \ S_{t}=\sum\limits_{i=1}^{t}b^{t}_{i}x_{i}x_{i}^{\top}\in\mathbb{R}^{d\times d},$ (1)
-for arbitrary coefficients $a^{t}_{i}\in\mathbb{R}$ and $b^{t}_{i}\in\mathbb{R}$. This formulation includes many practical schemes, see Table 1.
+for arbitrary coefficients $a^{t}_{i}\in\mathbb{R}$ and $b^{t}_{i}\in\mathbb{R}$.
+This formulation includes many practical schemes; see Table 1 in the appendix.
 To express this problem compactly, we rewrite it in matrix form.
@@
 We stack the data as rows into a matrix $X\in\mathbb{R}^{n\times d}$.
-Then, all terms of the sums in (1) can then be expressed compactly [^1] For simplicity, we slightly abuse the notation and make $S$ a matrix of shape $n\times d^{2}$ instead of a tensor of size $n\times d\times d$.
-as $Y=A_{1}X$ and $S=A_{2}(X\operatorname{\!\bullet\!}X)$, where the second matrix consists of the data vectors stacked as rows, and $\bullet$ denotes the Face-Splitting Product or transposed Khatri-Rao product (Esteve et al., 2009), $(X\operatorname{\!\bullet\!}X)_{ij}:=(x_{i}\otimes x_{i})_{j}=\textsf{vec}(x_{i}x_{i}^{\top})_{j}$, where $\otimes$ is the Kronecker product of two vectors.
-To protect the privacy of individual data elements, we rely on the notion of differential privacy, considering neighboring datasets as those differing by a single data point; see e.g.Dwork & Roth (2014) for an introduction.
-Recently, a series of works have developed effective methods for privately estimating individual matrices in the aforementioned product form by means of the the matrix mechanism (Li et al., 2015; Choquette-Choo et al., 2024; Denisov et al., 2022; Choquette-Choo et al., 2023b; Dvijotham et al., 2024; Kalinin & Lampert, 2024; Henzinger et al., 2024; Henzinger & Upadhyay, 2025; Fichtenberger et al., 2023; Henzinger et al., 2023; Choquette-Choo et al., 2023a; Kairouz et al., 2021; McMahan et al., 2024).
+Then, all terms of the sums in (1) can then be expressed compactly [^1] We slightly abuse the notation and make $S$ a matrix of shape $n\times d^{2}$ instead of a tensor of size $n\times d\times d$.
+as $Y=A_{1}X$ and $S=A_{2}(X\operatorname{\bullet}X)$, where the second matrix consists of the data vectors stacked as rows, and $\bullet$ denotes the Face-Splitting Product or transposed Khatri-Rao product (esteve2009, ), $(X\operatorname{\bullet}X)_{ij}:=(x_{i}\otimes x_{i})_{j}=\textsf{vec}(x_{i}x_{i}^{\top})_{j}$, where $\otimes$ is the Kronecker product of two vectors.
+To protect the privacy of individual data elements, we rely on the notion of differential privacy, considering neighboring datasets as those differing by a single data point; see e.g.(dwork2014algorithmic, ) for an introduction.
+Recently, a series of works have developed effective methods for privately estimating individual matrices in the aforementioned product form by means of the matrix mechanism (li2015matrix, ; CNPBINDPL, ; Denisov, ; MultiEpoch, ; dvijotham2024efficient, ; kalinin2024, ; henzinger2024, ).
 Its core insight is that, in the continual release setting, outputs at later steps should require less noise to be made private than earlier ones because more data points contributed to their computation.
@@
 This estimate is $(\epsilon,\delta)$-differentially private if the noise magnitude is at least $\text{sens}(CX)\cdot\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ denotes the variance required in the Gaussian distribution to ensure $(\epsilon,\delta)$-DP for sensitivity $1$ queries [^2] Here and in the following we do not provide a formula for $\sigma_{\epsilon,\delta}$, because closed-form expressions have been shown to be suboptimal in some regimes.
-Instead, we recommend determining its value numerically, such as in (Balle & Wang, 2018).
+Instead, we recommend determining its value numerically, such as in (balle2018improving, ).
 For reference, typical values of $\sigma_{\epsilon,\delta}$ lie between approximately $0.5$ (low privacy, e.g.
@@
 $\text{sens}(F)=\max_{X\sim X^{\prime}}\|F(X)\!-\!F(X^{\prime})\|_{\text{F}}$ (2)
-where $X\sim X^{\prime}$ denote any two data matrices, that are neighboring, i.e. identical except for one of the data vectors.
+where $X\sim X^{\prime}$ denote two neighboring data matrices, i.e. identical except for one of the data vectors.
 3 Joint Moment Estimation (JME)
-Algorithm 1 Joint Moment Estimation (JME)
-0:  stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$
-0:  workload matrices $A_{1},A_{2}\in\mathbb{R}^{n\times n}$
-0:  privacy parameters $(\epsilon,\delta)$
-0:  noise shaping matrices $C_{1},C_{2}\in\mathbb{R}^{n\times n}$ (invertible with decreasing column norm) default: $\text{I}_{n\times n},\text{I}_{n\times n}$
-  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-DP Gaussian mechanism
-  $\lambda\leftarrow\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$ // scaling parameter,   where $c_{1}=\frac{8}{11+5\sqrt{5}}$, and $c_{d}=2$ for $d\geq 2$
-  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // joint sensitivity
-  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
-  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d^{2}}$ // 2nd moment noise
-  for $t=1,2,\dots,n$ do
-     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
-     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
-     yield   $\widehat{Y_{t}}=\sum\limits_{i=1}^{t}[A_{1}]_{t,i}\widehat{x_{i}},\ \ \widehat{S_{t}}=\sum\limits_{i=1}^{t}[A_{2}]_{t,i}\widehat{x_{i}\otimes x_{i}}$
-  end for
 We now introduce our main algorithmic contribution: the Joint Moment Estimation (JME) algorithm for solving the problem of differentially private (weighted) moment estimation in a continual release setting.
@@
 Consequently, the fact that we also privately estimate the second moment does not increase the necessary noise level for the first moment, a property we call (second moment) privacy for free.
+Algorithm 1 Joint Moment Estimation (JME)
+stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$,
+workload matrices $A_{1},A_{2}\in\mathbb{R}^{n\times n}$
+privacy parameters $(\epsilon,\delta)$
+noise shaping matrices $C_{1},C_{2}\in\mathbb{R}^{n\times n}$ (invertible, decreasing column norm), defaults: Id
+$\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-DP Gaussian mechanism
+$\lambda\leftarrow\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$ with $c_{1}=\frac{8}{11+5\sqrt{5}}$, and $c_{d}=2$ for $d\geq 2$  // scaling parameter
+$s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // joint sensitivity
+$Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$, $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d^{2}}$ // 1st and 2nd moment noise
+for $t=1,2,\dots,n$ do
+     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$ and $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
+     yield   $\widehat{Y_{t}}=\sum\limits_{i=1}^{t}[A_{1}]_{t,i}\widehat{x_{i}},\ \ \widehat{S_{t}}=\sum\limits_{i=1}^{t}[A_{2}]_{t,i}\widehat{x_{i}\otimes x_{i}}$
+end for
 Note that privacy for free is a quite remarkable property of JME.
-Generally, when using the same data more than once for private computations, more noise for each of them is required to ensure the overall privacy, following, e.g., the composition theorems of DP (Kairouz et al., 2015).
+Generally, when using the same data more than once for private computations, more noise for each of them is required to ensure the overall privacy, following, e.g., the composition theorems of DP (kairouz15, ).
 In some situations, however, one might also prefer a more flexible way to trade-off between the two moment estimates.
 For this, we present $\lambda$-JME in the appendix (Algorithm 3), a variant of JME that has $\lambda$ as a free hyper-parameter.
-3.1 Properties of JME
+Properties of JME
 For some inputs $X=(x_{1},\dots,x_{n})\in\mathbb{R}^{n\times d}$, let $Y=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $S=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the matrices of their first and second moments after each step, and let $\widehat{Y}=(Y_{1},\dots,Y_{n})\in\mathbb{R}^{n\times d}$ and $\widehat{S}=(S_{1},\dots,S_{n})\in\mathbb{R}^{n\times d^{2}}$ be the private estimates computed by Algorithm 1 with workload matrix $A$ and noise shaping matrix $C$.
-Then, the following properties hold.
+Then, we have the following.
 Theorem 3.1 (Noise properties of JME).
@@
 Theorem 3.3 (Utility of JME).
-For any input $X$, the expected approximation error of Algorithm 1 for the first and second moments are
+For any input $X$, let $c_{d}$ be as defined in Algorithm 1. Then the expected approximation error of Algorithm 1 for the first and second moments are
 $\sqrt{\mathbb{E}\|Y-\widehat{Y}\|^{2}_{\text{F}}}$ $=2\zeta\sqrt{d}\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}\|A_{1}C_{1}^{-1}\|_{\text{F}}.$ (3)
-$\sqrt{\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}}$ $=2\zeta\sqrt{c_{d}}d\sigma_{\epsilon,\delta}\|C_{2}\|_{1\to 2}\|A_{2}C_{2}^{-1}\|_{\text{F}},$ (4)
-with $c_{d}$ as defined in Algorithm 1.
-3.2 Proofs
+$\sqrt{\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}}$ $=2\zeta^{2}\sqrt{c_{d}}d\sigma_{\epsilon,\delta}\|C_{2}\|_{1\to 2}\|A_{2}C_{2}^{-1}\|_{\text{F}}.$ (4)
 Proof of Theorem 3.1.
-The statement follows from the fact that for any $t=1,\dots,n$, JME’s estimates $\widehat{x_{t}}$ and $\widehat{x_{t}\otimes x_{t}}$ of $x_{t}$ and $x_{t}\otimes x_{t}$ are unbiased with symmetric noise distribution because they are constructed by adding zero-mean Gaussian noise.
+The statement follows from the fact that for any $t=1,\dots,n$, JME’s estimates $\widehat{x_{t}}$ of $x_{t}$ and $\widehat{x_{t}\otimes x_{t}}$ of $x_{t}\otimes x_{t}$ are unbiased with symmetric noise distribution because they are constructed by adding zero-mean Gaussian noise.
 ∎
-Sketch of the Proof of Theorem 3.2.
-By the properties of the Gaussian mechanism, it suffices to show that the overall sensitivity of jointly estimating both moments has at most the value $s=2\zeta\|C_{1}\|_{1\to 2}$, as stated in Algorithm 1.
+Proof sketch of Theorem 3.2.
+The result follows using Gaussian mechanism once we show that the overall sensitivity of jointly estimating both moments has at most $s=2\zeta\|C_{1}\|_{1\to 2}$, as stated in Algorithm 1.
 Definition 3.4.
@@
 $=\sup_{X\sim X^{\prime}}\!\!\Big{[}\|C_{1}(X\!-\!X^{\prime})\|_{\text{F}}^{2}+\lambda\|C_{2}(X\operatorname{\bullet}X\!-\!X^{\prime}\operatorname{\bullet}X^{\prime})\|_{\text{F}}^{2}\Big{]}.$
-The following Lemma (proved in the appendix) characterizes the values of the joint sensitivity as a function of $\lambda$.
+The following lemma (shown in the appendix) characterizes the values of the joint sensitivity as a function of $\lambda$.
 Lemma 3.5 (Joint Sensitivity).
-Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns. Then, for any $\lambda>0$ holds:
+Assume that the matrices $C_{1}$ and $C_{2}$ have norm-decreasing columns [^3] The decreasing column norm structure ensures that the optimum is simultaneously achieved for the first columns of the matrices $C_{1}$ and $C_{2}$.
+This can be relaxed by solving a maximization problem over the column indices, which could be done numerically.
+Moreover, the matrices that we have in mind for practical tasks are Toeplitz, for which the norm-decreasing property is fulfilled naturally..
+Then, for any $\lambda>0$ holds:
 $\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}\left(\frac{\lambda\zeta^{2}\|C_{2}\|_{1\to 2}^{2}}{\|C_{1}\|_{1\to 2}^{2}}\right),$ (6)
 where $\|\cdot\|_{1\to 2}$ denotes the maximum column norm, corresponding to the norm of the first column of $C_{1}$ and $C_{2}$, respectively. The function $r_{d}(\nu)$ is given by
-$r_{d}(\nu)=\begin{cases}\frac{1}{8}(3-\tau)^{2}(\nu\tau+1+\nu),&\text{if }\nu>\frac{11+5\sqrt{5}}{8},\,d=1,\\[5.0pt] 2+2\nu+\frac{1}{2\nu},&\text{if }\nu>\frac{1}{2},\,d>1,\\[5.0pt] 4,&\text{otherwise,}\end{cases}$ (7)
-where $\tau=\sqrt{1-\frac{2}{\nu}}$.
+$r_{d}(\nu)=\begin{cases}\frac{1}{8}(3-\tau)^{2}(\nu\tau+1+\nu),\text{ with }\tau=\sqrt{1-\frac{2}{\nu}},\quad&\text{if }\nu>\frac{11+5\sqrt{5}}{8},\,d=1,\\ 2+2\nu+\frac{1}{2\nu},&\text{if }\nu>\frac{1}{2},\,d>1,\\[5.0pt] 4,&\text{otherwise.}\end{cases}$ (7)
 Algorithm 1 uses the scaling parameter $\lambda=\|C_{1}\|_{1\to 2}^{2}/(c_{d}\zeta^{2}\|C_{2}\|_{1\to 2}^{2})$, so, by Lemma 3.5, the sensitivity of estimating both moment has the value $\sqrt{\zeta^{2}\|C_{1}\|_{1\to 2}^{2}r_{d}(c_{d})}$.
-The choice of $c_{d}$ implies that $r_{d}(c_{d})=4$, which establishes that $\text{sens}_{\lambda}(C_{1},C_{2})=s$, which concludes the proof of Theorem 3.2.
+The choice of $c_{d}$ implies that $r_{d}(c_{d})=4$, which yields $\text{sens}_{\lambda}(C_{1},C_{2})=s$.
+This concludes the proof of Theorem 3.2.
 ∎
-Sketch of the Proof of Theorem 3.3.
+Proof sketch of Theorem 3.3.
 The identities follow from the general properties of the matrix mechanism.
@@
 Consequently, $\widehat{Y}-Y=A_{1}C^{-1}_{1}Z_{1}$, and hence
-$\operatorname{\mathbb{E}}_{Z_{1}}\|\widehat{Y}-Y\|^{2}_{\text{F}}$ $=\|A_{1}C^{-1}_{1}\|^{2}_{\text{F}}\cdot\sigma^{2}_{\epsilon,\delta}s^{2}d$
+$\operatorname{\mathbb{E}}_{Z_{1}}\|\widehat{Y}-Y\|^{2}_{\text{F}}$ $=\|A_{1}C^{-1}_{1}\|^{2}_{\text{F}}\cdot\sigma^{2}_{\epsilon,\delta}s^{2}d$ (8)
 Equation (3) follows by inserting $s=2\zeta\|C_{1}\|_{1\to 2}$.
@@
 3.1 Comparison with alternative techniques
-Besides JME, a number of alternatives methods are possible that could be used to privately estimate the first and second moments.
-In this section, we introduce some of them and discuss their relation to JME.
-3.3.1 Independent Moment Estimation (IME)
-A straight-forward way to privately estimate first and second moments is to estimate and privatize both of them separately, where the necessary amount of noise is determined by the composition theorem of the Gaussian mechanism (Abadi et al., 2016) (see Algorithm 4 in the appendix).
+Besides JME, alternatives methods are possible that could be used to privately estimate the first and second moments. In this section, we introduce some of them and discuss their relation to JME.
+Independent Moment Estimation (IME).
+A straight-forward way to privately estimate first and second moments is to estimate and privatize both of them separately, where the necessary amount of noise is determined by the composition theorem of the Gaussian mechanism (abadi2016deep, ) (see Algorithm 4 in the appendix).
 IME resembles JME in the sense that (i) separate estimates of the moments are created, and (ii) the privatized results are unbiased estimators.
@@
 The proof can be found in the appendix. Figure 1 visualizes this property graphically.
-3.3.2 Concatenate-and-Split (CS)
+Concatenate-and-split (CS).
 In the special case where the two noise shaping matrices are meant to be the same (e.g.
@@
 Algorithm 5 in the appendix provides pseudocode for this concatenate-and-split (CS) method.
+Figure 1: Approximation errors for the 1st and 2nd moments across three methods (see text for details): $\lambda$-JME with $\lambda$, IME with varying $\alpha$ parameters, and CS with varying $\tau$ parameter.
+$\lambda$-JME Pareto-dominates the other two methods.
 When applicable, CS is easy to implement and produces unbiased moment estimates.
-However, like IME, it has an unfavorable privacy-accuracy trade-off curve compared to JME, as formalized in the following theorem.
-Figure 1: Approximation errors for the first and second moments ($d=10$, $n=100$, $C_{1}=C_{2}=I,A=E_{1}$, $\zeta=1$, and $\sigma=1$) across three methods: $\lambda$-JME with $\lambda$, IME with varying $\alpha$ parameters, and CS with varying $\tau$ parameter.
-One can see that $\lambda$-JME Pareto-dominates the other two methods.
+However, like IME, it has an unfavorable privacy-accuracy trade-off curve compared to JME, as formalized in the following theorem and shown in the appendix.
 Theorem 3.7 (JME vs CS).
 For any $\epsilon,\delta>0$, $\lambda$-JME Pareto-dominates CS with respect to the approximation error for the first vs second moment estimates.
-The proof can be found in the appendix. Figure 1 visualizes these cases (Theorems 3.6 and 3.7) graphically.
-3.3.3 Post-processing (PP)
+Figure 1 visualizes these cases (Theorems 3.6 and 3.7) graphically in an exemplary setting $(d=10,n=100,C_{1}=C_{2}=I,A=E_{1},\zeta=1,\sigma_{\epsilon,\delta}=1/2)$.
+Post-processing (PP).
 Post-processing (PP) is another easy-to-use method for joint moment estimation.
-It has appeared in the literature (Sheffet, 2019), at least in its naive form without the matrix factorization mechanism.
+It has appeared in the literature dp_covariance_wishart_dist , at least in its naive form without the matrix factorization mechanism.
 For any $x_{i}$, it first computes a private estimate $\widehat{x_{t}}$ by adding sufficient noise to it.
@@
 Algorithm 6 in the appendix provides pseudocode.
-(a) $d=1$, $C_{1}=C_{2}=\text{I}$
-(b) $d=1$, $C_{1}=C_{2}=A^{1/2}$
-(c) $d=1000$, $C_{1}=C_{2}=\text{I}$
-(d) $d=1000$, $C_{1}=C_{2}=A^{1/2}$
+(a) $d=1$, trivial
+(b) $d=1$, square root
+(c) $d=10^{3}$, trivial
+(d) $d=10^{3}$, square root
 Figure 2: Expected error of second moment estimation with JME versus PP with and without debiasing ($A=E_{1}$ (prefix sum), $n=1000$) In line with our analysis, for $d=1$ JME consistently achieves a higher quality than PP.
@@
 For given privacy parameters, PP and JME use the same noise strength toprivatize the first moment and therefore their estimates are of equal quality.
-For the second moment, PP and JME differ in their characteristics between the low privacy (small noise variance) and the high privacy (large noise variance) regime.
+For the second moment, their characteristics differ between the low privacy (small noise variance) and the high privacy (large noise variance) regime.
 To allow for a quantitative comparison, we derive characterizations of the approximation quality of PP.
@@
 Assume the same setting as for Theorem 3.3, except that we compute the estimates $\widehat{Y}$ and $\widehat{S}_{\text{PP}}$ with the PP method.
+Let $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}(Q)^{\top}$.
 Then, the expected approximation error of the second moment satisfies:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}=\overbrace{d\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}^{\text{bias}}$
-$\ \ +d(d+1)\sigma^{4}\zeta^{2}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (8)
-$\ \ +2(d+1)\sigma^{2}\zeta^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{\top}A_{2}\circ Q)XX^{T}),$
-where $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}(Q)^{\top}$. For debiased PP, the term marked ”bias” does not occur.
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}=16d(d+1)\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (9)
+$\quad\quad+8(d+1)\sigma_{\epsilon,\delta}^{2}\zeta^{2}\|C_{1}\|_{1\to 2}^{2}\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{\top}A_{2}\circ Q)XX^{T})\overbrace{+16d\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}^{\text{bias}}.$
+For debiased PP, the term marked "bias" does not occur.
 In order to get a better impression of the relation between PP and JME, we first study the special case of a trivial factorization.
 Corollary 3.9.
-Assume that JME and PP use trivial factorizations, i.e. $C_{1}=C_{2}=I$. Then, JME’s expected approximation error for the second moment is
-$\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}$ $=4c_{d}d^{2}\sigma^{2}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$ (9)
+Assume that JME and PP use trivial factorizations, i.e., $C_{1}=C_{2}=I$. Then, JME’s expected approximation error for the second moment is
+$\mathbb{E}\|S-\widehat{S}\|^{2}_{\text{F}}$ $=4c_{d}d^{2}\sigma_{\epsilon,\delta}^{2}\zeta^{4}\|A_{2}\|_{\text{F}}^{2}$ (10)
 and the corresponding value for PP is
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=\big{(}d(d+1)\sigma^{4}+2(d+1)\sigma^{2}\big{)}\zeta^{2}\|A_{2}\|_{\text{F}}^{2}$
-$\quad+\underbrace{d\sigma^{4}\zeta^{2}\|\sum_{i}(A_{2})_{[i,\cdot]}\|_{2}^{2}}_{\text{bias}},$ (10)
-where $\sum_{i}(A_{2})_{[i,\cdot]}$ is the row-wise summation of $A_{2}$.
-For debiased PP, the term marked ”bias” does not occur.
-Proof. The proofs follow directly from Theorem 3.3 and Lemma 3.8 by observing that $\|C_{1}\|_{1\to 2}=\|C_{2}\|_{1\to 2}=1$.
-In dimension $d=1$, JME has an error of $c_{1}\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ versus $(2\sigma^{4}+4\sigma^{2})\|A_{2}\|_{\text{F}}^{2}$ for (debiased) PP.
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=\big{(}16d(d+1)\zeta^{4}\sigma_{\epsilon,\delta}^{4}+8(d+1)\sigma_{\epsilon,\delta}^{2}\zeta^{2}\big{)}\|A_{2}\|_{\text{F}}^{2}\overbrace{+16d\sigma_{\epsilon,\delta}^{4}\zeta^{4}\|\sum_{i}(A_{2})_{[i,\cdot]}\|_{2}^{2}}^{\text{bias}},$
+where $\sum_{i}(A_{2})_{[i,\cdot]}$ is the row-wise summation of $A_{2}$. For debiased PP, the term marked "bias" does not occur.
+Proof.
+The proofs follow directly from Theorem 3.3 and Lemma 3.8 by observing that $\|C_{1}\|_{1\to 2}=\|C_{2}\|_{1\to 2}=1$.
+In dimension $d=1$, $\zeta=1$, JME has an error of $4c_{1}\sigma_{\epsilon,\delta}^{2}\|A_{2}\|_{\text{F}}^{2}$ versus $(32\sigma_{\epsilon,\delta}^{4}+16\sigma_{\epsilon,\delta}^{2})\|A_{2}\|_{\text{F}}^{2}$ for (debiased) PP.
 Since $c_{1}<1$, for one-dimensional data, JME’s error is always lower than PP’s (see Figure 2 for visual confirmation).
-In high dimensions ($d\gg 1$), the terms quadratic in $d$ dominate, so the comparison is between $2\sigma^{2}\|A_{2}\|_{\text{F}}^{2}$ for JME and $d^{2}\sigma^{4}\|A_{2}\|_{\text{F}}^{2}$ for PP.
-Consequently, at least for $\sigma\geq\sqrt{2}$, JME achieves privacy with less added noise than PP.
-In the regime of low privacy ($\sigma\to 0$) in high dimension $(d\gg 1)$, PP can be expected to result in higher accuracy estimates than JME, because the terms involving $\sigma^{4}$ make only minor contributions.
+In high dimensions ($d\gg 1$), the terms quadratic in $d$ dominate, so the comparison is between $8\sigma_{\epsilon,\delta}^{2}\|A_{2}\|_{\text{F}}^{2}$ for JME and $16d^{2}\sigma_{\epsilon,\delta}^{4}\|A_{2}\|_{\text{F}}^{2}$ for PP.
+Consequently, at least for $\sigma_{\epsilon,\delta}\geq\frac{1}{\sqrt{2}}$, JME achieves privacy with less added noise than PP.
+In the regime of low privacy ($\sigma_{\epsilon,\delta}\to 0$) in high dimension $(d\gg 1)$, PP can be expected to result in higher accuracy estimates than JME, because the terms involving $\sigma_{\epsilon,\delta}^{4}$ make only minor contributions.
 For settings with general noise shaping matrix, we cannot provide an exact comparison between PP and JME, because the $\sup_{X}$-term in Equation 9 has no closed-form solution.
-Instead, we derive upper and lower bounds (Lemma C.1 in the Appendix), and provide a numeric comparison in Figure 2.
-unbiased? symmetric noise? privacy-for-free?
-Method 1st / 2nd moment 1st / 2nd moment 2nd moment recommended regime of use
-JME (proposed) ✓/ ✓ ✓/ ✓ high privacy or low dimensionality
-IME ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
-CS ✓/ ✓ ✓/ ✓ never (Pareto-inferior to JME)
-PP ✓/✗ ✓/ ✗ low privacy in high dimensions
-PP (Debiased) ✓/✓ ✓/ ✗ low privacy in high dimensions
-Table 2: Overview of the properties for the methods discussed in Section 3.
-Properties unbiased and symmetric noise apply to each entry, stated separately for first and second moment.
-privacy-for-free applies only for the second moment.
+Instead, we derive upper and lower bounds (Lemma D.2), and provide a numeric comparison in Figure 2.
 4 Applications
@@
 The proofs for all theoretical results can be found in the appendix.
-4.1 Private Gaussian density estimation
+Private Gaussian density estimation
 The maximum likelihood solutions to Gaussian density estimation from i.i.d.
@@
 Using the alternative identity $\widehat{\Sigma}_{t}=\frac{1}{t}\sum_{i=1}^{t}x_{i}x^{\top}_{i}-\mu_{t}\mu^{\top}_{t}$, one sees that the task can indeed be solved in a continuous release setting and that only estimates of the first and second moments are required.
-(a) $d=1$, $n=1000$
-(b) $d=100$, $n=1000$
-Figure 3: Expected error of the estimated covariance matrix with JME versus PP (with trivial factorizations).
-For $d=1$, JME consistently achieves quality better than or on par with PP.
-For $d=100$, JME is preferable to PP only in the high privacy regime.
 To do so privately, we use the averaging workload matrix $V=(a^{t}_{i})$ with $a^{t}_{i}=\frac{1}{t}$ for $1\leq i\leq t$ and $a^{t}_{i}=0$ otherwise, and we compute private estimates $\widehat{\mu}:=\widehat{VX}$ (private mean) and $\widehat{\Sigma}=\widehat{V(X\operatorname{\bullet}X)}-\widehat{(VX)}\operatorname{\bullet}\widehat{(VX)}$ (private covariance) using JME.
@@
 The following Theorem establishes the approximation quality of the covariance estimate.
+In this and the following sections we denote $\sigma=2\sigma_{\epsilon,\delta}$, referring to the strength of the noise we add to the first moment for $\zeta=1$ and trivial factorization.
 Theorem 4.1 (Private covariance matrix estimation with JME).
@@
 Then it holds:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=(c_{d}d^{2}+2d+2)\sigma^{2}H_{n,1}$ (11)
-$\quad+d(d+1)\sigma^{4}H_{n,2},$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=(c_{d}d^{2}+2d+2)\sigma^{2}H_{n,1}+d(d+1)\sigma^{4}H_{n,2},$ (11)
 with $c_{d}$ as in Theorem 3.3, and $H_{n,m}:=\sum\limits_{k=1}^{n}\frac{1}{k^{m}}$.
@@
 Let $\widehat{\Sigma}_{\text{PP}}$ be the result of the above construction, where privacy is obtained by running PP with noise strength $\sigma$ and debiasing.
+Let $S(n,d,\sigma):=d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}$.
 Then, for the expected error of the covariance matrix estimate it holds:
-$d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}$
-$\qquad+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
-$\leq\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ (12)
-$\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}.$
+$S(n,d,\sigma)-2(d+1)\sigma^{2}H_{n,3}\leq\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}\leq S(n,d,\sigma).$ (12)
 Comparison.
@@
 Figure 6 shows a numerical plot that confirms this observation.
-Note that our results match the lower bounds established by Kamath et al.
-(2020), who proved that private covariance estimation in the Frobenius norm requires $\Omega(d^{2})$ samples.
-4.2 Private Adam optimization
-The Adam optimizer (Diederik & Ba, 2015), has become a de-facto standard for optimization in deep learning.
+Note that our results match the lower bounds established by G.
+Kamath 2020 (dp_covariance_lower_bound, ), who proved that private covariance estimation in the Frobenius norm requires $\Omega(d^{2})$ samples.
+Private Adam optimization
+The Adam optimizer (kingma2014adam, ), has become a de-facto standard for optimization in deep learning.
 The defining property of Adam is its update rule, $\theta_{i}\leftarrow\theta_{i-1}-\alpha m_{i}/(\sqrt{v_{i}}+\epsilon)$, where $\alpha$ is a learning rate, and $m_{i}$ and $v_{i}$ are exponentially running averages of computed model gradients and componentwise squared model gradients, respectively.
 In the context of our work, these quantities correspond to a weighted first-moment vector and the diagonal of the weighted second-moment matrix.
-Previous attempts to make Adam differentially private relied on postprocessing (Anil et al., 2022; Li et al., 2022), potentially with debiasing (Tang et al., 2024), i.e.
+Previous attempts to make Adam differentially private relied on postprocessing (anil2021large, ; li2022large, ), potentially with debiasing (dp_adam_is_not_adam, ), i.e.
 they privatized the model gradients and derived the squared values from these.
@@
 $\mathbb{E}_{Z}\|\widehat{D}-D\|^{2}$ $=2d\sigma^{2}\cdot\|A_{2}\|_{\text{F}}^{2},$ (13)
-$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=(2d\sigma^{4}+4\sigma^{2})\cdot\|A_{2}\|_{\text{F}}^{2}$
-$\quad+\underbrace{d\sigma^{4}\cdot\big{\|}\sum_{i}(A_{2})_{i}\big{\|}_{2}^{2}}_{\text{bias}},$ (14)
-where $A_{2}$ is the workload matrix obtained from the coefficients of Adam’s exponentially weighted averaging operations. The term marked ”bias” disappears if PP is debiased.
-The proof of the first part follows directly from Theorem 3.3. The error for PP is computed separately in Lemma C.4 in the appendix.
+$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=(2d\sigma^{4}+4\sigma^{2})\cdot\|A_{2}\|_{\text{F}}^{2}\ \overbrace{+d\sigma^{4}\cdot\big{\|}\sum_{i}(A_{2})_{i}\big{\|}_{2}^{2}}^{\text{bias}},$ (14)
+where $A_{2}$ is the workload matrix obtained from the coefficients of Adam’s exponentially weighted averaging operations. The term marked "bias" disappears if PP is debiased.
+The proof of the first part follows directly from Theorem 3.3. The error for PP is computed separately in Lemma D.5 in the appendix.
 This lemma shows that as long as $\sigma\geq 1$, the error introduced by JME is strictly lower than that of PP (i.e.
-classic DP-Adam (Li et al., 2022)) and even the debiased version of PP (i.e.
-DP-AdamBC (Tang et al., 2024)).
+classic DP-Adam (li2022large, )) and even the debiased version of PP (i.e.
+DP-AdamBC (dp_adam_is_not_adam, )).
 Figure 5 illustrates the relation in a prototypical case.
-(a) trivial factorization
-(b) square root factorization
-Figure 4: Expected error of the estimated covariance vector for Adam with JME versus PP ($d=10^{6}$; $n=1000$). JME provides a better estimate in the mid to high privacy regime.
 5 Experiments
@@
 One can see that in the high-privacy regime (here: $\sigma=2$), on average, JME achieves a better estimate of the true density than PP, with and without debiasing.
-Figure 5: Approximation quality of JME and PP for private Gaussian density estimation ($d=5,n=100$) as average and standard deviation over 1000 runs.
-In the high-privacy regime ($\sigma=2$), JME achieves lower KL divergence than PP after $\approx 10$ samples.
+(a) $d=5,n=100,\sigma=2$
+(b) $d=10,n=200,\sigma=4$
+Figure 3: Approximation quality of JME and PP for private Gaussian density estimation as average and standard deviation over 1000 runs.
+In the high-privacy regime ($\sigma=2$ and $\sigma=4$), JME achieves lower KL divergence than PP after $\approx 10$ samples.
 (a) batchsize $B\!=\!256$, $\varepsilon\approx 1.7$
@@
 We train a convolutional network on the CIFAR-10 dataset with DP-Adam, which is privatized either with JME or PP.
-Because gradients could have arbitrarily large number, we apply gradient clipping to a model-selected threshold in both methods.
+Because gradients could have arbitrarily large norm, we apply gradient clipping to a model-selected threshold in both methods.
 In addition, we include a heuristic baseline that uses the concatenate-and-split techniques in which not the norm of the gradient vector but the norm of the concatenation vector is clipped.
@@
 6 Related works
-The problem of differentially private moments (and the related problem of covariance estimation) has a rich history of development, with optimal results known in the central model of privacy (Smith, 2011) as well as the local model of privacy (Duchi et al., 2013), and also in the worst-case setting.
-Sheffet (2019) proposes three differentially private algorithms for approximating the second-moment matrix, each ensuring positive definiteness.
-The related setting of covariance estimation has been studied in the worst-case setting when the data comes from a bounded $\ell_{2}$ ball by several works for approximate differential privacy (Achlioptas & McSherry, 2007; Blum et al., 2005; Dwork et al., 2014; Mangoubi & Vishnoi, 2022, 2023; Upadhyay, 2018) with the works of Amin et al.
-(2019) and Kapralov & Talwar (2013) presenting a pure differentially private algorithm.
-Covariance estimation is also used as a subroutine in mean-estimation work under various distributional assumptions (Kamath, 2024).
+The problem of differentially private moments (and the related problem of covariance estimation) has a rich history of development, with optimal results known in the central model of privacy (smith2011privacy, ) as well as the local model of privacy (duchi2013local, ), and also in the worst-case setting.
+Sheffet (dp_covariance_wishart_dist, ) proposes three differentially private algorithms for approximating the second-moment matrix, each ensuring positive definiteness.
+The related setting of covariance estimation has been studied in the worst-case setting when the data comes from a bounded $\ell_{2}$ ball by several works for approximate differential privacy (achlioptas2007fast, ; blum2005practical, ; dwork2014analyze, ; mangoubi2022re, ; mangoubi2023private, ; upadhyay2018price, ) with the works of Amin et al.
+(dp_covariance, ) and Kapralov & Talwar (kapralov2013differentially, ) presenting a pure differentially private algorithm.
+Covariance estimation is also used as a subroutine in mean-estimation work under various distributional assumptions (kamath2024broader, ).
 However, none of these approaches are directly applicable to the continual release model and they offer improvements over the Gaussian mechanism only in a very high privacy regime.
+Precise sensitivity analysis, as a technique to improve differential privacy guarantees, has also been intensively used in the literature (lebeda2024correlated, ; wilkins2024exact, ; joseph2024privately, ; lebeda2025better, ).
+The matrix mechanism has also gained a lot of attention recently due to its application of continually releasing prefix sums in private online optimization BandedMatrix ; Kairouz ; efficient_streaming ; henzinger2025improved ; fichtenberger2023constant ; henzinger2023almost ; kalinin2025back ; henzinger2025binned .
 7 Summary and Discussion
@@
 We presented the Joint Moment Estimation (JME) method, which solves the problem by exploiting the recently proposed matrix mechanism with carefully tuned noise level.
-As a result, JME produces unbiased estimates of both moments while requiring less noise to be added than baseline methods, at least is the high privacy regime.
+As a result, JME produces unbiased estimates of both moments while requiring less noise to be added than baseline methods, at least in the high privacy regime.
 We applied JME to private Gaussian density estimation and model training with Adam, demonstrating improved performance in high-privacy regimes both theoretically and practically.
 Despite the promising results, several open question remain.
-In particular, would like to explore if postprocessing is indeed the optimal strategy in a low-privacy regime, or if a better privacy-utility trade-off is still possible.
+In particular, we would like to explore if postprocessing is indeed the optimal strategy in a low-privacy regime, or if a better privacy-utility trade-off is still possible.
 Furthermore, we plan to explore the possibility of problem-specific factorizations, which could be fused with the proposed method.
@@
 We are also grateful to Mher Safaryan for his contributions to discussions on DP-Adam.
-Additionally, we thank Ryan McKenna for suggesting Joint Clipping (CS) as a baseline.
+Additionally, we thank Ryan McKenna for suggesting Joint Clipping as a baseline.
 Jalaj Upadhyay’s research was funded by the Rutgers Decanal Grant no.
@@
 A.1 Proof of Theorem 3.2 and Lemma 3.5
-The privacy of Algorithm 1 follows from the properties of matrix mechanism (Li et al., 2015), with a precise estimate of the sensitivity of the joint estimation process that we will introduce and discuss later in this section.
+The privacy of Algorithm 1 follows from the properties of matrix mechanism [34], with a precise estimate of the sensitivity of the joint estimation process that we will introduce and discuss later in this section.
 By means of the matrix factorization mechanism with $A_{1}=B_{1}C_{1}$ and $A_{2}=B_{2}C_{2}$, we write the joint moment estimate as,
@@
 where $\lambda>0$ is an arbitrary trade-off parameter that we will adjust optimally later.
-To make the estimate private, we privatize the rightmost matrix in (15), which contains the data, by adding suitable scaled Gaussian noise.
+To make the estimate private, we privatize the rightmost matrix in (15), which contains the data, by adding suitably scaled Gaussian noise.
 The subsequent matrix multiplication then acts on private data, so its result is also private.
 It remains to show that the noise level specified in Algorithm 1 suffices to guarantee $(\epsilon,\delta)$-privacy.
-For this, we denote by $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ be the set of input sequences where each $d$-dimensional vectors has bounded $\ell_{2}$ norm.
+For this, we denote by $\mathcal{X}=\{(x_{1},\dots,x_{n})\;:\;\max_{i}\|x_{i}\|_{2}\leq\zeta\wedge x_{i}\in\mathbb{R}^{d}\}$ be the set of input sequences where each $d$-dimensional vector has bounded $\ell_{2}$ norm.
 The sensitivity of a computation is the maximal amount by which the result differs between two input sequences $X,X^{\prime}\in\mathcal{X}$, which are identical except for a single data vector at an arbitrary index (denoted by $X\sim X^{\prime}$).
@@
 A straightforward calculation shows
-$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\max_{i=1,\dots,n}\sup_{\|x\|\leq\zeta,\|y\|\leq\zeta}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2}$
-$\;=\frac{\zeta^{2}}{\alpha^{2}_{1}}\left[\sup_{\|x\|\leq 1,\|y\|\leq 1}\|x-y\|_{2}^{2}+\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\|x\otimes x-y\otimes y\|_{2}^{2}\right]$
+$\text{sens}_{\lambda}^{2}(C_{1},C_{2})$ $=\max_{i=1,\dots,n}\sup_{\|x\|\leq\zeta,\|y\|\leq\zeta}\alpha^{2}_{i}\|x-y\|_{2}^{2}+\lambda\beta^{2}_{i}\|x\otimes x-y\otimes y\|_{2}^{2}$ (22)
+$\;=\zeta^{2}\alpha^{2}_{1}\left[\sup_{\|x\|\leq 1,\|y\|\leq 1}\|x-y\|_{2}^{2}+\frac{\lambda\zeta^{2}\beta^{2}_{1}}{\alpha^{2}_{1}}\|x\otimes x-y\otimes y\|_{2}^{2}\right]$
 $\;=\zeta^{2}\alpha^{2}_{1}r_{d}(\lambda\zeta^{2}\beta^{2}_{1}/\alpha^{2}_{1})$
 completing the proof of Corollary A.3. ∎
-Corollary 1 implies that the joint estimate (15) will be $(\epsilon,\delta)$-private, if we use noise of strength at least, $\sigma=\zeta^{2}\alpha_{1}r_{d}(\lambda\zeta^{2}\beta_{1}/\alpha_{1})\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ is the noise strength required for the Gaussian mechanism with sensitivity $1$.
+Corollary A.3 implies that the joint estimate (15) will be $(\epsilon,\delta)$-private, if we use noise of strength at least, $\sigma=\zeta^{2}\alpha_{1}r_{d}(\lambda\zeta^{2}\beta_{1}/\alpha_{1})\sigma_{\epsilon,\delta}$, where $\sigma_{\epsilon,\delta}$ is the noise strength required for the Gaussian mechanism with sensitivity $1$.
 This finishes the proof of the Lemma 3.5.
@@
 $\mathbb{E}_{Z}\|Y-\hat{Y}\|^{2}_{\text{F}}$ $=\mathbb{E}_{Z}\|B_{1}Z_{1}\|^{2}=d\sigma^{2}\|B_{1}\|^{2}_{\text{F}}.$ (25)
-Inserting $B_{1}=A_{1}C_{1}^{-1}$ and $\sigma=2\zeta\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}$, Equation (3) follows. Analogoulsy, for $Z_{2}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d\times d}$,
+Inserting $B_{1}=A_{1}C_{1}^{-1}$ and $\sigma=2\zeta\sigma_{\epsilon,\delta}\|C_{1}\|_{1\to 2}$, Equation (3) follows. Analogously, for $Z_{2}\sim[\mathcal{N}(0,\sigma^{2}\text{I})]^{d\times d}$,
 $\mathbb{E}_{Z}\|S-\hat{S}\|^{2}_{\text{F}}$ $=\mathbb{E}_{Z}\|\frac{1}{\sqrt{\lambda^{*}}}B_{2}Z_{2}\|^{2}=\frac{d^{2}\sigma^{2}}{\lambda^{*}}\|B_{2}\|^{2}_{\text{F}}.$ (26)
@@
 Specifically:
-$r^{\text{diag}}_{d}(\lambda)=\sup\limits_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\!\!\!\big{[}\|x-y\|_{2}^{2}+\lambda\|\operatorname{\operatorname{diag}}(x\otimes x)-\operatorname{\operatorname{diag}}(y\otimes y)\|_{2}^{2}\big{]}=r_{d}(\lambda)$
+$r^{\text{diag}}_{d}(\lambda)=\sup\limits_{\|x\|_{2}\leq 1,\|y\|_{2}\leq 1}\!\!\!\big{[}\|x-y\|_{2}^{2}+\lambda\|\operatorname{\operatorname{diag}}(x\otimes x)-\operatorname{\operatorname{diag}}(y\otimes y)\|_{2}^{2}\big{]}=r_{d}(\lambda)$ (27)
 For simplicity, we denote $x\circ x=\operatorname{\operatorname{diag}}(x\otimes x)$.
@@
 For any vectors $x,y\in\mathbb{R}^{d}$, where $d\geq 3$, there exist vectors $x^{\prime},y^{\prime}\in\mathbb{R}^{d-1}$ that for any $\lambda>0$ satisfies the inequality:
-$\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\leq\|x^{\prime}-y^{\prime}\|_{2}^{2}+\lambda\|x^{\prime}\circ x^{\prime}-y^{\prime}\circ y^{\prime}\|_{2}^{2}.$
+$\|x-y\|_{2}^{2}+\lambda\|x\circ x-y\circ y\|_{2}^{2}\leq\|x^{\prime}-y^{\prime}\|_{2}^{2}+\lambda\|x^{\prime}\circ x^{\prime}-y^{\prime}\circ y^{\prime}\|_{2}^{2}.$ (29)
 We apply this lemma recursively to prove that $r^{\text{diag}}_{d}(\lambda)=r^{\text{diag}}_{2}(\lambda)$ for all $d\geq 2$. The proof of the lemma can be found later in the appendix.
@@
 Appendix B Additional materials
+Table 1: Common workload matrices for moment estimation in (1).
+method symbol coefficients
+prefix sum $E_{1}$ $\mathbbm{1}{\{i\leq t\}}$
+exponential (weight $\beta$) $E_{\beta}$ $\beta^{t-i}\mathbbm{1}\{i\leq t\}$
+standard average $V$ $\frac{1}{t}\mathbbm{1}\{i\leq t\}$
+sliding window (size $k$) $W_{k}$ $\frac{1}{k}\mathbbm{1}\{t-k\,<\,i\,\leq\,t\}$
+(a) trivial factorization
+(b) square root factorization
+Figure 5: Expected error of the estimated covariance vector for Adam with JME versus PP ($d=10^{6}$; $n=1000$). JME provides a better estimate in the mid to high privacy regime.
+(a) $d=1$, $n=1000$
+(b) $d=100$, $n=1000$
+Figure 6: Expected error of the estimated covariance matrix with JME versus PP (with trivial factorizations).
+For $d=1$, JME consistently achieves quality better than or on par with PP.
+For $d=100$, JME is preferable to PP only in the high privacy regime.
 (a) $d=100,\beta=0.9$, $C_{1}=C_{2}=I$
@@
 In line with our analysis, incorporating matrix factorization significantly improves the quality of both methods, particularly in the high privacy regime.
+Table 2: CIFAR-10 experiments with two different privacy budgets, $\varepsilon\approx 1.7$ and $\varepsilon\approx 0.16$ for $\delta=10^{-6}$, for four methods: DP-Adam with and without debiasing, JME, and Joint Clipping.
+The average and standard deviation errors are based on 3 runs.
 Method Epoch 1 Epoch 2 Epoch 3 Epoch 4 Epoch 5 Epoch 6 Epoch 7 Epoch 8 Epoch 9 Epoch 10
@@
 DP-Adam 26.58 $\pm$ 0.43 31.45 $\pm$ 0.43 35.02 $\pm$ 1.16 36.39 $\pm$ 0.58 38.15 $\pm$ 0.95 39.31 $\pm$ 0.52 40.33 $\pm$ 0.33 41.47 $\pm$ 1.05 42.42 $\pm$ 0.94 43.07 $\pm$ 0.86
-Table 3: CIFAR-10 experiments with two different privacy budgets, $\varepsilon\approx 1.7$ and $\varepsilon\approx 0.16$ for $\delta=10^{-6}$, for four methods: DP-Adam with and without debiasing, JME, and Joint Clipping.
-The average and standard deviation errors are based on 3 runs.
+Table 3: Hyperparameters for CIFAR-10 Experiments.
+Medium-privacy experiments use a batch size of $256$, compared to $1$ in the high-privacy regime, while using a noise multiplier of $\sigma_{\varepsilon,\delta}=1$ for the medium-privacy regime and $\sigma_{\varepsilon,\delta}=2$ for the high-privacy regime.
+JME and joint clipping require an additional hyperparameter—scaling—which is optimized to find the best value for those runs.
+We also find it helpful to clip the updates; for this, we use the same clipping norm.
 Noise Mult Batch Size Method lr Scaling ($\lambda,\tau)$ eps Clip Norm
@@
 DP-Adam $10^{-7}$ $10^{-8}$ 1
-Table 4: Hyperparameters for CIFAR-10 Experiments.
-Medium-privacy experiments use a batch size of $256$, compared to $1$ in the high-privacy regime, while using a noise multiplier of $\sigma_{\varepsilon,\delta}=1$ for the medium-privacy regime and $\sigma_{\varepsilon,\delta}=2$ for the high-privacy regime.
-JME and joint clipping require an additional hyperparameter—scaling—which is optimized to find the best value for those runs.
-We also find it helpful to clip the updates; for this, we use the same clipping norm.
-Definition B.1 (Face-Splitting Product).
+Appendix C Algorithms
+Algorithm 2 Differentially Private JME Adam
+Input: Initial model $\theta_{0}\in\mathbb{R}^{d}$, dataset $D$, batchsize $b$, matrices $C_{\beta_{1}},C_{\beta_{2}}\in\mathbb{R}^{n\times n}$, model loss $\ell(\theta,d)$, clipnorm $\zeta$, noise multiplier $\sigma_{\epsilon,\delta}\geq 0$, learning rate $\alpha>0$, and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\varepsilon=10^{-8}$.
+$m_{0}\leftarrow 0$   // first moment initialization.
+$v_{0}\leftarrow 0$   // second moment initialization.
+$\lambda,s_{\lambda}=\cdot\textit{Joint-sens}(C_{\beta_{1}},C_{\beta_{2}})$   // joint sensitivity
+$Z_{1},Z_{2}\sim N(0,\sigma_{\epsilon,\delta}^{2}s_{\lambda}^{2}I_{d})$   // noise generating
+for $i=1,2,\dots,n$ do
+     $S_{i}\leftarrow\{d_{1},\dots,d_{m}\}\subseteq D$ select a data batch
+     $g_{j}\leftarrow\nabla_{\theta}\ell(\theta_{i-1},d_{j}))\quad\text{for $j=1,\dots,m$}$
+     $x_{i}\leftarrow\sum_{j=1}^{m}\min(1,\zeta/||g_{j}||)g_{j}$
+     $\widehat{x}_{i}\leftarrow x_{i}+\zeta[C_{\beta_{1}}^{-1}Z_{1}]_{[i,\cdot]}$
+     $\widehat{x^{2}}_{i}\leftarrow x_{i}^{2}+\lambda^{-1/2}\zeta[C_{\beta_{2}}^{-1}Z_{2}]_{[i,\cdot]}$
+     $m_{i}\leftarrow m_{i-1}\beta_{1}+(1-\beta_{1})\widehat{x}_{i}$
+     $v_{i}\leftarrow v_{i-1}\beta_{2}+(1-\beta_{2})\widehat{x^{2}}_{i}$
+     $\widehat{m}_{i}=m_{i}/(1-\beta_{1}^{i})$   // bias-correction
+     $\widehat{v}_{i}=v_{i}/(1-\beta_{2}^{i})$   // bias-correction
+     $\theta_{i}\leftarrow\theta_{i-1}-\alpha\widehat{m}_{i}/(\sqrt{\widehat{v}_{i}}+\epsilon)$
+end for
+$\Theta=(\theta_{1},\dots,\theta_{n})$
+Algorithm 3 $\lambda$-JME
+input stream vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
+workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
+noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms)  (default: $\text{I}_{n\times n}$)
+privacy parameters $(\epsilon,\delta)$
+$\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
+$s\leftarrow\zeta\|C_{1}\|_{1\to 2}r_{d}\left(\frac{\lambda\zeta^{2}\|C_{2}\|_{1\to 2}^{2}}{\|C_{1}\|_{1\to 2}^{2}}\right)^{1/2}$ // joint sensitivity
+$Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
+$Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
+for $t=1,2,\dots,n$ do
+     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
+     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
+[…]
+end for
+Algorithm 4 $\alpha$-IME (IME with budget split parameter $\alpha\in(0,1)$)
+input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
+workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
+noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
+privacy parameters $(\epsilon,\delta)$
+$\sigma\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
+privacy trade-off $\alpha\in(0,1)$
+$\sigma_{1}\leftarrow\frac{\sigma}{\sqrt{\alpha}}$
+$s_{1}\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
+$Z_{1}\sim\big{[}\mathcal{N}(0,\sigma^{2}_{1}s^{2}_{1})\big{]}^{n\times d}$ // 1st moment noise
+$\sigma_{2}\leftarrow\frac{\sigma}{\sqrt{1-\alpha}}$
+$s_{2}\leftarrow\sqrt{2}\zeta^{2}\|C_{2}\|_{1\to 2}$ // sensitivity of 2nd moment
+$Z_{2}\sim\big{[}\mathcal{N}(0,\sigma^{2}s_{2}^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
+for $t=1,2,\dots,n$ do
+     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
+     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
+[…]
+end for
+Algorithm 5 $\tau$-CS (CS with 2nd moment rescaling parameter $\tau>0$)
+input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with
+workload matrices $A=(a^{t}_{k})\in\mathbb{R}^{n\times n}$
+input dimension $d$, bound on input norm $\zeta>0$
+privacy parameters $(\epsilon,\delta)$
+(optional) noise shaping matrix $C$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
+$\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
+$s\leftarrow 2\zeta\sqrt{1+\tau\zeta^{2}}$ // sensitivity based on norm of concatenated data
+$Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d(d+1)}$ // noise matrix for concatenated data
+for $t=1,2,\dots,n$ do
+     $\widetilde{x_{t}}=\big{(}\,x_{t},\sqrt{\tau}\textsf{vec}(x_{t}\otimes x_{t})\,\big{)}$
+     $\widehat{\widetilde{x_{t}}}\leftarrow\widetilde{x_{t}}+[C^{-1}Z]_{[t,\cdot]}$
+[…]
+end for
+Algorithm 6 PP
+input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
+workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
+noise shaping matrix $C_{1}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
+privacy parameters $(\epsilon,\delta)$
+$\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
+$s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
+$Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
+for $t=1,2,\dots,n$ do
+     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z]_{[t,\cdot]}$
+     $b_{t}\leftarrow I_{d\times d}\times\sigma_{\epsilon,\delta}^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{k=1}^{n}(A_{2})_{t,k}(C_{1}C_{1}^{\top})^{-1}_{k,k}$ // bias term (optional)
+     $\widehat{x_{t}\otimes x_{t}}\leftarrow\widehat{x_{t}}\otimes\widehat{x_{t}}-b_{t}$
+[…]
+end for
+Appendix D Technical Proofs
+Definition D.1 (Face-Splitting Product).
 Let $A=(a_{i}^{\top})_{i=1}^{n}$ with $a_{i}\in\mathbb{R}^{d_{1}}$ and $B=(b_{i}^{\top})_{i=1}^{n}$ with $b_{i}\in\mathbb{R}^{d_{2}}$.
@@
   • Associativity $(A\operatorname{\bullet}B)\operatorname{\bullet}C=A\operatorname{\bullet}(B\operatorname{\bullet}C)$.
-  • Frobenious norm $\|A\operatorname{\!\bullet\!}B\|_{\text{F}}=\|B\operatorname{\!\bullet\!}A\|_{\text{F}}$, even so $A\operatorname{\!\bullet\!}B\neq B\operatorname{\!\bullet\!}A$.
-Appendix C Technical Proofs
-C.1 Bounds on the expected approximation error for PP with non-trivial factorization
-Lemma C.1.
+  • Frobenius norm $\|A\operatorname{\bullet}B\|_{\text{F}}=\|B\operatorname{\bullet}A\|_{\text{F}}$, even though $A\operatorname{\bullet}B\neq B\operatorname{\bullet}A$.
+D.1 Bounds on the expected approximation error for PP with non-trivial factorization
+Lemma D.2.
 $\|A_{2}C_{1}^{-1}\|^{2}_{\text{F}}$ $\leq\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}\big{(}(A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top})XX^{\top}\big{)}\leq\sum_{ij}\big{|}\,[A_{2}^{\top}A_{2}\circ C_{1}^{-1}C_{1}^{-\top}]_{i,j}\,\big{|}$ (31)
-Proof of Lemma C.1.
+Proof of Lemma D.2.
 The proof is elementary.
@@
 Let $\beta=\langle x,y\rangle$, where $-1\leq\beta\leq 1$.
-The expression $-\beta-2\lambda\beta^{2}$ is maximized at $\beta=-\frac{1}{2\lambda}$ when $\lambda\geq\frac{1}{2}$, or at $\beta=-1$ when $\lambda<\frac{1}{2}$.
+The expression $-2\beta-2\lambda\beta^{2}$ is maximized at $\beta=-\frac{1}{2\lambda}$ when $\lambda\geq\frac{1}{2}$, or at $\beta=-1$ when $\lambda<\frac{1}{2}$.
 If $\beta=-1$, then $x=-y$ and the objective function becomes $4$.
@@
 We aim to evaluate the expected squared Frobenius norm of the error:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(X+C_{1}^{-1}Z_{1}))-A_{2}(X\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}$
-$\quad=2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{\mathcal{S}_{1}}+2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\!\bullet\!}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\!\bullet\!}X)\rangle_{\text{F}}}_{\mathcal{S}_{2}}+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{\mathcal{S}_{3}}.$
-We compute those terms separately.
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\operatorname{\bullet}(X+C_{1}^{-1}Z_{1}))-A_{2}(X\operatorname{\bullet}X)\|_{\text{F}}^{2}$ (52)
+$\quad=2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\operatorname{\bullet}C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{\mathcal{S}_{1}}+2\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\langle A_{2}(X\operatorname{\bullet}C_{1}^{-1}Z_{1}),A_{2}(C_{1}^{-1}Z_{1}\operatorname{\bullet}X)\rangle_{\text{F}}}_{\mathcal{S}_{2}}$
+$\qquad\qquad+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{\mathcal{S}_{3}}.$
+We compute those terms separately. We add component-wise independent Gaussian noise $Z_{1}\sim\mathcal{N}(0,\sigma^{2})$ to the first moment, where $\sigma=2\zeta\sigma_{\epsilon,\delta}$.
 Step 1. Bound $\mathcal{S}_{1}$.
@@
 We now deal with two cases. When $i=j$, we compute:
-$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}&=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}\\ &\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (52)
+$\begin{split}\mathbb{E}&(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}\\ &=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}\\ &=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}\\ &\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (56)
 On the other hand, when $i\neq j$, then
-$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}&=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\right)^{2}=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (53)
+$\begin{split}\mathbb{E}(C_{1}^{-1}Z_{1})_{t_{1},i}(C_{1}^{-1}Z_{1})_{t_{1},j}(C_{1}^{-1}Z_{1})_{t_{2},i}(C_{1}^{-1}Z_{1})_{t_{2},j}&=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\right)^{2}\\ &=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}Q_{t_{1},t_{2}}^{2}.\end{split}$ (57)
 Using equation (56) and equation (57) in equation (55), we get
-$\mathcal{S}_{3}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
+$\mathcal{S}_{3}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ (58)
+$=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
 $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+(d+1)Q_{t_{1},t_{2}}^{2})$
@@
 Adding equation (53) to equation (55), we obtain:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+2(d+1)\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|S-\widehat{S}_{\text{PP}}\|^{2}_{\text{F}}$ $=d(d+1)\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (59)
+$\quad+2(d+1)\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
 $\quad+d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
@@
 $[\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))]_{k,i,j}$ $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,i}\right)\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)$ (60)
-$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
+$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}$
+$=\sigma^{2}\delta_{i,j}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}(A_{2})_{k,t}Q_{t,t}.$
 The Frobenius norm of this bias is:
-$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
+$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$ (61)
 $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
 If we subtract this bias from the estimate, it will increase the error by the aforementioned quantity due to the Frobenius norm of the bias but will decrease the error by two scalar products with the $A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))$ term:
-$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\!\bullet\!}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
-Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction. ∎
+$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\operatorname{\bullet}(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$ (62)
+Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction.
+Substituting back $\sigma=2\zeta\sigma_{\epsilon,\delta}$, we obtain the proposed equality.
+∎
 We collect some useful proposition that would be useful for our analysis.
-Proposition C.2.
+Proposition D.3.
 Let $V$ and $X$ be a given fixed matrix and $Z$ be a Gaussian matrix of appropriate dimension. Then
-$\mathbb{E}_{Z}\|VZ\operatorname{\!\bullet\!}VZ\|_{\text{F}}^{2}=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.$
+$\mathbb{E}_{Z}\|VZ\operatorname{\bullet}VZ\|_{\text{F}}^{2}=d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.$ (63)
 Proof.
@@
 This completes the proof of the proposition. ∎
-Proposition C.3.
+Proposition D.4.
 Let $V$ and $X$ be a given fixed matrix and $Z$ be a Gaussian matrix of appropriate dimension. Then
-$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\!\bullet\!}VX\|_{\text{F}}^{2}.$
+$\mathbb{E}_{Z}\langle(VZ)\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ)\rangle=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\bullet}VX\|_{\text{F}}^{2}.$ (65)
 Proof.
 The result follows using the following calculation:
-$\mathbb{E}_{Z}\langle(VZ)\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ)\rangle$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ_{1})_{k,i}(VX)_{k,j}(VZ_{1})_{k,j}(VX)_{k,i}=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\left(\sum\limits_{i=1}^{d}(VZ_{1})_{k,i}(VX)_{k,i}\right)^{2}$
+$\mathbb{E}_{Z}\langle(VZ)\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ)\rangle$ $=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\sum\limits_{i,j}^{d}(VZ_{1})_{k,i}(VX)_{k,j}(VZ_{1})_{k,j}(VX)_{k,i}$ (66)
+$=\mathbb{E}_{Z}\sum\limits_{k=1}^{n}\left(\sum\limits_{i=1}^{d}(VZ_{1})_{k,i}(VX)_{k,i}\right)^{2}$
 $=\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}=\frac{1}{d}\mathbb{E}_{Z}\|VZ_{1}\operatorname{\bullet}VX\|_{\text{F}}^{2}$
@@
 We first recall that, if $Z\sim\mathcal{N}(\mu,\Sigma)$, then for any matrix $A$, we have $AZ\sim\mathcal{N}(A\mu,A\Sigma A^{\top})$. This implies that, when $\Sigma=\mathbb{I}$ and $\mu=0$, we have
-$\mathbb{E}_{Z}\|AZ\|_{\text{F}}^{2}=\|A\|_{\text{F}}^{2}.$
+$\mathbb{E}_{Z}\|AZ\|_{\text{F}}^{2}=\|A\|_{\text{F}}^{2}.$ (67)
 Recall that $\widehat{\mu}=V(X+Z_{1})$ is a running mean and $\widehat{\Sigma}=V(X\operatorname{\bullet}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))$ is a running covariance matrix, with independent noise $Z_{1},Z_{2}\in\mathcal{N}(0,\sigma^{2})^{n\times d^{2}}$, the clipping norm $\zeta=1$, and $\lambda=\lambda^{*}=c_{d}^{-1}$ as defined in equation (23).
 Using the associativity of Face-splitting product and the Pythogorean theorem, we have
-$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}$
+$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\bullet}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)\|_{\text{F}}^{2}$ (68)
 $\quad=\mathbb{E}_{Z_{1},Z_{2}}\|\lambda^{-1/2}VZ_{2}-(VZ_{1})\operatorname{\bullet}(VX)-(VX)\operatorname{\bullet}(VZ_{1})-(VZ_{1})\operatorname{\bullet}(VZ_{1}))\|_{\text{F}}^{2}$
-$\quad=\mathbb{E}_{Z_{2}}\|\lambda^{-1/2}VZ_{2}\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle$
-$\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}}+{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}}+2{\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}$
-Using Proposition C.2 and Proposition C.3, we therefore have
+$\quad=\mathbb{E}_{Z_{2}}\|\lambda^{-1/2}VZ_{2}\|_{\text{F}}^{2}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$
+$\qquad\qquad+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle$
+$\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}}+{\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}}$
+$\qquad\qquad+2{\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle}$
+Using Proposition D.3 and Proposition D.4, we therefore have
 $\begin{split}&\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\bullet}X+\lambda^{-1/2}Z_{2})-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)\|_{\text{F}}^{2}\\ &\quad=c_{d}\sigma^{2}d^{2}\|V\|_{\text{F}}^{2}+2\left(1+{1\over d}\right){\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}.\end{split}$ (69)
@@
 Then the unbiased approximation error is
-$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\!\bullet\!}X+Z_{2})-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)+\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$
+$\mathbb{E}_{Z_{1},Z_{2}}\|V(X\operatorname{\bullet}X+Z_{2})$ $-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)$ (76)
+$+\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$
 We have already computed the first three terms inside the expectation. We now compute the bias term error:
-$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
+$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$ (77)
 Bias reduction procedure decreases the expected error of (74) by
-$-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
+$-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\rangle+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$ (78)
+$=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}=d\sigma^{4}H_{n,2}.$
 Resulting in the final error of:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=\sigma^{2}(c_{d}d^{2}+2d+2)\sum\limits_{k=1}^{n}\frac{1}{k}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}\|_{\text{F}}^{2}$ $=\sigma^{2}(c_{d}d^{2}+2d+2)\sum\limits_{k=1}^{n}\frac{1}{k}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$ (79)
 $=\sigma^{2}(c_{d}d^{2}+2d+2)H_{n,1}+d(d+1)\sigma^{4}H_{n,2}.$
@@
 Let us denote the covariance matrix estimated via Post-Processing (PP) without bias correction as $\widehat{\Sigma}^{b}_{\text{PP}}$. Then, the approximation error has the following form:
-$\begin{split}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}=&\mathbb{E}_{Z_{1}}\|V((X+Z_{1})\operatorname{\!\bullet\!}(X+Z_{1}))-(V(X+Z_{1})\operatorname{\!\bullet\!}V(X+Z_{1}))-V(X\operatorname{\!\bullet\!}X)+(VX)\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}\\ &\quad=\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})+V(Z_{1}\operatorname{\!\bullet\!}X)+V(X\operatorname{\!\bullet\!}Z_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VX)-(VX)\operatorname{\!\bullet\!}(VZ_{1})-(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}))\|_{\text{F}}^{2}\\ &\quad=\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}}_{A_{1}}-2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}Z_{1}),(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{2}}+2\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\!\bullet\!}X)\|_{\text{F}}^{2}}_{A_{3}}\\ &\hskip 56.9055pt+2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),V(X\operatorname{\!\bullet\!}Z_{1})\rangle}_{A_{4}}-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle}_{A_{5}}\\ &\hskip 56.9055pt-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\!\bullet\!}X),(VZ_{1})\operatorname{\!\bullet\!}(VX)\rangle}_{A_{6}}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VX)\|_{\text{F}}^{2}+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\!\bullet\!}(VX),(VX)\operatorname{\!\bullet\!}(VZ_{1})\rangle\end{split}$ (63)
+$\begin{split}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}=&\mathbb{E}_{Z_{1}}\|V((X+Z_{1})\operatorname{\bullet}(X+Z_{1}))-(V(X+Z_{1})\operatorname{\bullet}V(X+Z_{1}))-V(X\operatorname{\bullet}X)+(VX)\operatorname{\bullet}(VX)\|_{\text{F}}^{2}\\ &\quad=\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}}_{A_{1}}-2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}Z_{1}),(VZ_{1})\operatorname{\bullet}(VZ_{1})\rangle}_{A_{2}}+2\underbrace{\mathbb{E}_{Z_{1}}\|V(Z_{1}\operatorname{\bullet}X)\|_{\text{F}}^{2}}_{A_{3}}\\ &\hskip 56.9055pt+2\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),V(X\operatorname{\bullet}Z_{1})\rangle}_{A_{4}}-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),(VX)\operatorname{\bullet}(VZ_{1})\rangle}_{A_{5}}\\ &\hskip 56.9055pt-4\underbrace{\mathbb{E}_{Z_{1}}\langle V(Z_{1}\operatorname{\bullet}X),(VZ_{1})\operatorname{\bullet}(VX)\rangle}_{A_{6}}+2\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VX)\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+\mathbb{E}_{Z_{1}}\|(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}\\ &\hskip 56.9055pt+2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VX),(VX)\operatorname{\bullet}(VZ_{1})\rangle\end{split}$ (80)
 The last three terms in the above expression evaluates to equation (LABEL:eq:simplifiedend). Therefore, in what follows, we bound $A_{1}$ to $A_{6}$.
@@
 Plugging equation (81) to equation (86) in equation (80), we get
-$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
+$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}+d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$ (87)
 $\quad-2d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)\left(\sum\limits_{t=1}^{n}V^{2}_{k,t}\right)$
-$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
+$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}$
+$\quad-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
 $\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$
 Recalling that the matrix $V$ is the averaging workload matrix $V=(a^{t}_{i})$ with $a^{t}_{i}=\frac{1}{t}$ for $1\leq i\leq t$ and $a^{t}_{i}=0$ otherwise, we get
-$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}-2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
+$\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}-2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$ (88)
 $\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{k}\frac{1}{k^{2}}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{k}\frac{1}{k^{2}}X_{t,j}\right)\left(\sum\limits_{t=1}^{k}\frac{1}{k}X_{t,j}\right)$
 $\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{k}\frac{1}{k}X_{t,j}\right)^{2}+d(d+2)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
-$=d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\underbrace{\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{1},:}\rangle\right]}_{T_{n}(X)}$
+$=d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
+$\quad+2(d+1)\sigma^{2}\underbrace{\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{1},:}\rangle\right]}_{T_{n}(X)}$
 Since every term except $T_{n}(X)$ is independent of $X$, to compute both upper and lower bounds on $\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$, it suffices to bound the supremum of the inner difference over $X$:
-$T_{n}:=\sup_{X\in\mathcal{X}}T_{n}(X)\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\underbrace{\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}}_{=\left\|\sum_{t}X_{t}\right\|^{2}_{2}\geq 0}\rangle\right].$
+$T_{n}:=\sup_{X\in\mathcal{X}}T_{n}(X)\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}\left[\sum\limits_{t=1}^{k}\langle X_{t,:},X_{t,:}\rangle-\frac{1}{k}\underbrace{\sum\limits_{t_{1},t_{2}}^{k}\langle X_{t_{1},:},X_{t_{2},:}}_{=\left\|\sum_{t}X_{t}\right\|^{2}_{2}\geq 0}\rangle\right].$ (89)
 Since the double sum is non-negative, this leads to a trivial upper bound $T_{n}\leq\sum\limits_{k=1}^{n}\frac{1}{k}=H_{n,1}$.
@@
 Therefore,
-$H_{n,1}-H_{n,3}\leq T_{n}\leq H_{n,1}.$
+$H_{n,1}-H_{n,3}\leq T_{n}\leq H_{n,1}.$ (90)
 Therefore, we have the following bounds for the approximation error without a bias correction:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}$
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}-2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k^{3}},$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}$ (91)
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}^{b}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d-1)\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}+dn\sigma^{4}-d^{2}\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k}-2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\frac{1}{k^{3}},$ (92)
 Now we will determine the bias term; let $\delta_{i=j}$ denote the Dirac-delta function, then the bias of $\widehat{\Sigma}_{\text{PP}}$ is
@@
 We also have the following set of equalities when $Z_{1}$ is a Gaussian matrix.
-$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$
-$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}=dn\sigma^{4}$
-$\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle_{\text{F}}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$
+$\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}$ (94)
+$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}$ $=\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}\right)^{2}=dn\sigma^{4}$ (95)
+$\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\rangle_{\text{F}}$ $=d\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}\right)=d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}$ (96)
 To remove the bias from the error we would need to add the following terms:
-$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\|_{\text{F}}^{2}+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1})\|_{\text{F}}^{2}-2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\!\bullet\!}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\!\bullet\!}Z_{1})\rangle$
+$\|\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\|_{\text{F}}^{2}+\|\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\|_{\text{F}}^{2}-2\langle\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\rangle$ (97)
 $\quad-2\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}(VZ_{1})\operatorname{\bullet}(VZ_{1})\rangle_{\text{F}}+4\mathbb{E}_{Z_{1}}\langle(VZ_{1})\operatorname{\bullet}(VZ_{1}),\mathbb{E}_{Z_{1}}V(Z_{1}\operatorname{\bullet}Z_{1})\rangle$
@@
 $=-dn\sigma^{4}-d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k^{2}}+2d\sigma^{4}\sum\limits_{k=1}^{n}\frac{1}{k}=-dn\sigma^{4}-d\sigma^{4}H_{n,2}+2d\sigma^{4}H_{n,1}.$
-Combining everything together we obtain:
-$\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$
-$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
+Combining everything together, we obtain:
+$\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $=d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}-2d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{3}$ (98)
+$\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}X_{t,j}^{2}$
+$\quad-4(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}X_{t,j}\right)\left(\sum\limits_{t=1}^{n}V_{k,t}X_{t,j}\right)$
 $\quad+2(d+1)\sigma^{2}\sum\limits_{k=1}^{n}\sum\limits_{t=1}^{n}V_{k,t}^{2}\sum\limits_{j=1}^{d}(VX)_{k,j}^{2}+d(d+1)\sigma^{4}\sum\limits_{k=1}^{n}\left(\sum\limits_{t=1}^{n}V_{k,t}^{2}\right)^{2}$
 Therefore,
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}\quad\text{and}$
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\leq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}\quad\text{and}$ (99)
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|\Sigma-\widehat{\Sigma}_{\text{PP}}\|_{\text{F}}^{2}$ $\geq d(d+1)\sigma^{4}H_{n,1}-d(d+1)\sigma^{4}H_{n,2}+2(d+1)\sigma^{2}H_{n,1}-2(d+1)\sigma^{2}H_{n,3},$ (100)
 which concludes the proof. ∎
@@
 Consider the difference in the objective function and $f(x_{,}i,y_{i})$ defined below:
-$f(x_{i},y_{i})$ $:=\left(\sqrt{x_{i}^{2}+x_{j}^{2}}+\sqrt{y_{i}^{2}+y_{j}^{2}}\right)^{2}+\lambda\left(x_{i}^{2}+x_{j}^{2}-y_{i}^{2}-y_{j}^{2}\right)^{2}$
-$g(x_{i},y_{i})$ $:=(x_{i}-y_{i})^{2}+(x_{j}-y_{j})^{2}+\lambda(x_{i}^{2}-y_{i}^{2})^{2}+\lambda(x_{j}^{2}-y_{j}^{2})^{2}.$
+$f(x_{i},y_{i})$ $:=\left(\sqrt{x_{i}^{2}+x_{j}^{2}}+\sqrt{y_{i}^{2}+y_{j}^{2}}\right)^{2}+\lambda\left(x_{i}^{2}+x_{j}^{2}-y_{i}^{2}-y_{j}^{2}\right)^{2}$ (101)
+$g(x_{i},y_{i})$ $:=(x_{i}-y_{i})^{2}+(x_{j}-y_{j})^{2}+\lambda(x_{i}^{2}-y_{i}^{2})^{2}+\lambda(x_{j}^{2}-y_{j}^{2})^{2}.$ (102)
 Note that $f(x_{i},y_{i})$ is the objective function. Now, after algebraic manipulation, we get
@@
 This concludes the proof. ∎
-Lemma C.4 (Expected Second Moment Error with PP).
+Lemma D.5 (Expected Second Moment Error with PP).
 Given the private estimation of the first moment $A_{1}(X+C_{1}^{-1}Z_{1})$, where $A_{1}=B_{1}C_{1}$, and the second moment $A_{2}(X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1})$, where $A_{2}=B_{2}C_{2}$, with independent noise $Z_{1}\in\mathcal{N}(0,\|C_{1}\|_{1\to 2}^{2}\sigma^{2})^{n\times d}$, the clipping norm $\zeta=1$, $d>1$, the expected squared Frobenius norm of the estimation error for the second moment satisfies:
 $\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}:$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$ (128)
-$=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
-$\quad+\underbrace{d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}_{\text{bias}}$
+$=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$
+$\qquad+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{T})$
+$\qquad+\underbrace{d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})}_{\text{bias}}$
 where $Q=C_{1}^{-1}C_{1}^{-\top}$ and $E_{Q}=\operatorname{\operatorname{diag}}(Q)\operatorname{\operatorname{diag}}^{\top}(Q)$.
@@
 We aim to evaluate the expected squared Frobenius norm of the error:
-$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$
+$\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}((X+C_{1}^{-1}Z_{1})\circ(X+C_{1}^{-1}Z_{1}))-A_{2}(X\circ X)\|_{\text{F}}^{2}$ (129)
 $\quad=4\underbrace{\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}}_{S_{1}}+\underbrace{\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}}_{S_{2}}$
 We compute those terms separately.
-$S_{1}$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$
+$S_{1}$ $=\sup_{X\in\mathcal{X}}\mathbb{E}\|A_{2}(X\circ C_{1}^{-1}Z_{1})\|_{\text{F}}^{2}=\sup_{X\in\mathcal{X}}\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}X_{t,j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$ (130)
 $=\sup_{X\in\mathcal{X}}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{1}}X_{t_{1},j}X_{t_{2},j}\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t_{1},r}(C_{1}^{-1})_{t_{2},r}\mathbb{E}(Z_{1})_{r,j}^{2}$
@@
 where $Q=C_{1}^{-1}C_{1}^{-T}$.
-$S_{2}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})^{2}_{t,j}\right)^{2}$
+$S_{2}$ $=\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\left(\sum\limits_{t=1}^{n}(A_{2})_{k,t}(C_{1}^{-1}Z_{1})^{2}_{t,j}\right)^{2}$ (131)
 $=\mathbb{E}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}.$
 First, we compute:
-$\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}$ $=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
+$\mathbb{E}(C_{1}^{-1}Z_{1})^{2}_{t_{1},j}(C_{1}^{-1}Z_{1})^{2}_{t_{2},j}$ $=\mathbb{E}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}(Z_{1})_{r,j}^{4}$ (132)
+$\quad+\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
 $\quad+2\mathbb{E}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}(Z_{1})_{r_{1},j}^{2}(Z_{1})_{r_{2},j}^{2}$
-$=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}$
+$=3\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r=1}^{n}(C_{1}^{-1})^{2}_{t_{1},r}(C_{1}^{-1})^{2}_{t_{2},r}$
+$\quad+\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})^{2}_{t_{1},r_{1}}(C_{1}^{-1})^{2}_{t_{2},r_{2}}$
 $\quad+2\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{r_{1}\neq r_{2}}^{n}(C_{1}^{-1})_{t_{1},r_{1}}(C_{1}^{-1})_{t_{1},r_{2}}(C_{1}^{-1})_{t_{2},r_{1}}(C_{1}^{-1})_{t_{2},r_{2}}$
@@
 Plugging it back, we get
-$\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$
+$\mathbb{E}\|A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$ (133)
 $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}(Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}+2Q_{t_{1},t_{2}}^{2})$
@@
 Adding these terms together we obtain:
-$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{\top})$
+$\sup_{X\in\mathcal{X}}\mathbb{E}_{Z}\|\widehat{D}_{\text{PP}}-D\|^{2}$ $=2d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}(Q\circ Q))$ (134)
+$\quad+4\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\cdot\sup_{X\in\mathcal{X}}\operatorname{\operatorname{tr}}((A_{2}^{T}A_{2}\circ Q)XX^{\top})$
 $\quad+d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\cdot\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
@@
 The expectation of $A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))]_{k,j}$ introduces a bias:
-$[\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))]_{k,j}$ $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$
+k,j $=\mathbb{E}\sum\limits_{t=1}^{n}(A_{2})_{k,t}\left(\sum\limits_{r=1}^{n}(C_{1}^{-1})_{t,r}(Z_{1})_{r,j}\right)^{2}$ (135)
 $=\sigma^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{t=1}^{n}\sum\limits_{r=1}^{n}(A_{2})_{k,t}(C_{1}^{-1})^{2}_{t,r}$
@@
 The Frobenius norm of this bias is:
-$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$
+$\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}$ $=\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\sum\limits_{k=1}^{n}\sum\limits_{j=1}^{d}\sum\limits_{t_{1},t_{2}}^{n}(A_{2})_{k,t_{1}}(A_{2})_{k,t_{2}}Q_{t_{1},t_{1}}Q_{t_{2},t_{2}}$ (136)
 $=d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$
 If we subtract this bias from the estimate, it will increase the error by the aforementioned quantity due to the Frobenius norm of the bias but will decrease the error by two scalar products with the $A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))$ term:
-$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$
+$\mathbb{E}\langle A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1})),\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\rangle=\|\mathbb{E}A_{2}((C_{1}^{-1}Z_{1})\circ(C_{1}^{-1}Z_{1}))\|_{\text{F}}^{2}.$ (137)
 Thus, we can eliminate the last term ($d\sigma^{4}\|C_{1}\|_{1\to 2}^{4}\operatorname{\operatorname{tr}}(A_{2}^{\top}A_{2}E_{Q})$) in the error sum via bias correction. ∎
-Appendix D Algorithms
-Algorithm 2 Differentially Private JME Adam
-  Input: Initial model $\theta_{0}\in\mathbb{R}^{d}$, dataset $D$, batchsize $b$, matrices $C_{\beta_{1}},C_{\beta_{2}}\in\mathbb{R}^{n\times n}$, model loss $\ell(\theta,d)$, clipnorm $\zeta$, noise multiplier $\sigma_{\epsilon,\delta}\geq 0$, learning rate $\alpha>0$, and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\varepsilon=10^{-8}$.
-  $m_{0}\leftarrow 0$   // first moment initialization.
-  $v_{0}\leftarrow 0$   // second moment initialization.
-  $\lambda,s_{\lambda}=\cdot\textit{Joint-sens}(C_{\beta_{1}},C_{\beta_{2}})$   // joint sensitivity
-  $Z_{1},Z_{2}\sim N(0,\sigma_{\epsilon,\delta}^{2}s_{\lambda}^{2}I_{d})$   // noise generating
-  for $i=1,2,\dots,n$ do
-     $S_{i}\leftarrow\{d_{1},\dots,d_{m}\}\subseteq D$ select a data batch
-     $g_{j}\leftarrow\nabla_{\theta}\ell(\theta_{i-1},d_{j}))\quad\text{for $j=1,\dots,m$}$
-     $x_{i}\leftarrow\sum_{j=1}^{m}\min(1,\zeta/||g_{j}||)g_{j}$
-     $\widehat{x}_{i}\leftarrow x_{i}+\zeta[C_{\beta_{1}}^{-1}Z_{1}]_{[i,\cdot]}$
-     $\widehat{x^{2}}_{i}\leftarrow x_{i}^{2}+\lambda^{-1/2}\zeta[C_{\beta_{2}}^{-1}Z_{2}]_{[i,\cdot]}$
-     $m_{i}\leftarrow m_{i-1}\beta_{1}+(1-\beta_{1})\widehat{x}_{i}$
-     $v_{i}\leftarrow v_{i-1}\beta_{2}+(1-\beta_{2})\widehat{x^{2}}_{i}$
-     $\widehat{m}_{i}=m_{i}/(1-\beta_{1}^{i})$   // bias-correction
-     $\widehat{v}_{i}=v_{i}/(1-\beta_{2}^{i})$   // bias-correction
-     $\theta_{i}\leftarrow\theta_{i-1}-\alpha\widehat{m}_{i}/(\sqrt{\widehat{v}_{i}}+\epsilon)$
-  end for
-  $\Theta=(\theta_{1},\dots,\theta_{n})$
-Algorithm 3 $\lambda$-JME
-0:  input stream vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
-0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
-0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms)  (default: $\text{I}_{n\times n}$)
-0:  privacy parameters $(\epsilon,\delta)$
-  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
-  $s\leftarrow\zeta\|C_{1}\|_{1\to 2}r_{d}\left(\frac{\lambda\zeta^{2}\|C_{2}\|_{1\to 2}^{2}}{\|C_{1}\|_{1\to 2}^{2}}\right)^{1/2}$ // joint sensitivity
-  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
-  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
-  for $t=1,2,\dots,n$ do
-     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
-     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+\lambda^{-1/2}[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
-[…]
-  end for
-Algorithm 4 $\alpha$-IME (IME with budget split parameter $\alpha\in(0,1)$)
-0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
-0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
-0:  noise shaping matrices $C_{1},C_{2}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
-0:  privacy parameters $(\epsilon,\delta)$
-0:  privacy trade-off $\alpha\in(0,1)$
-  $(\epsilon_{1},\delta_{1})\leftarrow(\alpha\epsilon,\alpha\delta)$ // 1st moment private budget
-  $(\epsilon_{2},\delta_{2})\leftarrow((1-\alpha)\epsilon,(1-\alpha)\delta)$ // 2nd moment private budget
-  $\sigma_{1}\leftarrow$ noise strength for $(\epsilon_{1},\delta_{1})$-dp Gaussian mechanism
-  $s_{1}\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
-  $Z_{1}\sim\big{[}\mathcal{N}(0,\sigma^{2}_{1}s^{2}_{1})\big{]}^{n\times d}$ // 1st moment noise
-  $\sigma_{2}\leftarrow$ noise strength for $(\epsilon_{2},\delta_{2})$-dp Gaussian mechanism
-  $s_{2}\leftarrow\sqrt{2}\zeta^{2}\|C_{2}\|_{1\to 2}$ // sensitivity of 2nd moment
-  $Z_{2}\sim\big{[}\mathcal{N}(0,\sigma^{2}s_{2}^{2})\big{]}^{n\times d\times d}$ // 2nd moment noise
-  for $t=1,2,\dots,n$ do
-     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z_{1}]_{[t,\cdot]}$
-     $\widehat{x_{t}\otimes x_{t}}\leftarrow x_{t}\otimes x_{t}+[C_{2}^{-1}Z_{2}]_{[t,\cdot,\cdot]}$
-[…]
-  end for
-Algorithm 5 $\tau$-CS (CS with 2nd moment rescaling parameter $\tau>0$)
-0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with
-0:  workload matrices $A=(a^{t}_{k})\in\mathbb{R}^{n\times n}$
-0:  input dimension $d$, bound on input norm $\zeta>0$
-0:  privacy parameters $(\epsilon,\delta)$
-0:  (optional) noise shaping matrix $C$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
-  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
-  $s\leftarrow 2\zeta\sqrt{1+\tau\zeta^{2}}$ // sensitivity based on norm of concatenated data
-  $Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d(d+1)}$ // noise matrix for concatenated data
-  for $t=1,2,\dots,n$ do
-     $\widetilde{x_{t}}=\big{(}\,x_{t},\sqrt{\tau}\textsf{vec}(x_{t}\otimes x_{t})\,\big{)}$
-     $\widehat{\widetilde{x_{t}}}\leftarrow\widetilde{x_{t}}+[C^{-1}Z]_{[t,\cdot]}$
-[…]
-  end for
-Algorithm 6 PP
-0:  input stream of vectors $x_{1},\dots,x_{n}\in\mathbb{R}^{d}$ with $\|x_{t}\|_{2}\leq\zeta$ for $\zeta>0$
-0:  workload matrices $A_{1}=(a^{t}_{k}),A_{2}=(b^{t}_{k})\in\mathbb{R}^{n\times n}$
-0:  noise shaping matrix $C_{1}$ (lower triangular, invertible, decreasing column norms) (default: $\text{I}_{n\times n}$)
-0:  privacy parameters $(\epsilon,\delta)$
-  $\sigma_{\epsilon,\delta}\leftarrow$ noise strength for $(\epsilon,\delta)$-dp Gaussian mechanism
-  $s\leftarrow 2\zeta\|C_{1}\|_{1\to 2}$ // sensitivity of 1st moment
-  $Z\sim\big{[}\mathcal{N}(0,\sigma_{\epsilon,\delta}^{2}s^{2})\big{]}^{n\times d}$ // 1st moment noise
-  for $t=1,2,\dots,n$ do
-     $\widehat{x_{t}}\leftarrow x_{t}+[C_{1}^{-1}Z]_{[t,\cdot]}$
-     $b_{t}\leftarrow I_{d\times d}\times\sigma_{\epsilon,\delta}^{2}\|C_{1}\|_{1\to 2}^{2}\sum\limits_{k=1}^{n}(A_{2})_{t,k}(C_{1}C_{1}^{\top})^{-1}_{k,k}$ // bias term (optional)
-     $\widehat{x_{t}\otimes x_{t}}\leftarrow\widehat{x_{t}}\otimes\widehat{x_{t}}-b_{t}$
-[…]
-  end for